### v5

# Algebroid Mappings and Their Equidistribution Theory

Abstract.
In this paper, the concept of algebroid mappings of complex manifolds is introduced based on that a large number of complex systems of PDEs admit multi-valued solutions that can be defined by a system of independent algebraic equations over the field of meromorphic functions. It is well-known that Nevanlinna theory is an important tool in complex ODE theory. To develop a similar tool applied to the study of algebroid solutions of complex systems of PDEs, one explores the equidistribution theory of algebroid mappings. Via uniformizating an algebroid mapping, we obtain a second main theorem of algebroid mappings from a complete Kähler manifold into a complex projective manifold provided that some certain conditions are imposed.
Key words and phrases:
Algebroid mappings; equidistribution; second main theorem; Kähler manifolds.
2010 Mathematics Subject Classification:
32H30; 32H25
1. Introduction
1.1. Motivation
Algebroid solutions of complex ODEs have been studied extensively since the notion of algebroid functions was introduced by H. Poincaré. It has been discovered that many complex ODEs have algebroid solutions. For instance, $4zW^{\prime 2}-1=0$ is of an algebroid solution $W=\sqrt{z}.$ What about an algebroid solution of a complex system of PDEs? Let $F_{i}(z,W_{1},\cdots,W_{d})$ $(i=1,\cdots,k)$ be functions of $W_{1},\cdots,W_{d}$ and their partial derivatives on $\mathbb{C}^{m}$ over the field $\mathscr{M}(\mathbb{C}^{m})$ of all meromorphic functions on $\mathbb{C}^{m},$ where $z=(z_{1},\cdots,z_{m})\in\mathbb{C}^{m}.$ A solution $(W_{1},\cdots,W_{n})$ of a given complex system of PDEs:
(1) $F_{i}\big(z,W_{1},\cdots,W_{d}\big)=0,\ \ \ \ i=1,\cdots,k$
is called an algebroid solution, if it is defined by a system of $d$ independent algebraic equations
(2) $P_{i}\big(z,W_{1},\cdots,W_{d}\big)=0,\ \ \ \ i=1,\cdots,d$
in $W_{1},\cdots,W_{d}$ over $\mathscr{M}(\mathbb{C}^{m}).$ Here, the “independent” means that
$\left|\frac{\partial(P_{1},\cdots,P_{d})}{\partial(W_{1},\cdots,W_{d})}\right|\not\equiv 0.$
Note that a meromorphic solution of Eq. (1) is always an algebroid solution. To our knowledge, a large number of complex systems of PDEs, particularly, lots of complex systems of algebraic PDEs, have non-meromorphic algebroid solutions.
Examples
  • The complex system of PDEs
    $\begin{cases}\frac{\partial W}{\partial z_{1}}=\frac{z_{2}}{2W}\\ \frac{\partial W}{\partial z_{2}}=\frac{z_{1}}{2W}\\ \frac{\partial^{2}W}{\partial z_{1}\partial z_{2}}=\frac{1}{4W}\end{cases}$
    on $\mathbb{C}^{2}$ has an algebroid solution
    $W=\sqrt{z_{1}z_{2}}.$
  • The complex system of PDEs
    $\begin{cases}W_{1}\frac{\partial^{2}W_{1}}{\partial z_{1}\partial z_{2}}+W_{2}\frac{\partial^{2}W_{2}}{\partial z_{1}\partial z_{2}}+\frac{\partial W_{1}}{\partial z_{1}}\frac{\partial W_{1}}{\partial z_{2}}+\frac{\partial W_{2}}{\partial z_{1}}\frac{\partial W_{2}}{\partial z_{2}}-\frac{1}{2}=0\\ W_{2}\frac{\partial^{2}W_{1}}{\partial z_{1}\partial z_{2}}+W_{1}\frac{\partial^{2}W_{2}}{\partial z_{1}\partial z_{2}}+\frac{\partial W_{1}}{\partial z_{1}}\frac{\partial W_{2}}{\partial z_{2}}+\frac{\partial W_{2}}{\partial z_{1}}\frac{\partial W_{1}}{\partial z_{2}}-1=0\end{cases}$
    on $\mathbb{C}^{2}$ has an algebroid solution $(W_{1},W_{2})$ defined by
    $W_{1}^{2}+W_{2}^{2}=z_{1}z_{2},\ \ \ \ W_{1}W_{2}=z_{1}+z_{2}.$
  • The complex system of PDEs
    $\begin{cases}\frac{\partial^{2}W_{1}}{\partial z_{1}\partial z_{2}}=\frac{1}{W_{2}}\\ \frac{\partial^{2}W_{2}}{\partial z_{1}\partial z_{2}}=\frac{1}{W_{1}}\end{cases}$
    on $\mathbb{C}^{2}$ has algebroid solutions $(W_{1},W_{2})$ of the form
    $W_{1}=W_{2}=\pm 2\sqrt{(\lambda z_{1}+c_{1})(\lambda^{-1}z_{2}+c_{2})},$
    where $\lambda\not=0,c_{1},c_{2}$ are arbitrary constants.
In the viewpoint of mappings, any algebroid solution $(W_{1},\cdots,W_{d})$ of Eq. (1) corresponds to a multi-valued analytic mapping
$F=\big[1:W_{1}:\cdots:W_{d}\big]:\ \mathbb{C}^{m}\to\mathbb{P}^{d}(\mathbb{C}).$
On the other hand, if the above $F$ is a multi-valued mapping defined by Eq. (2), then $(W_{1},\cdots,W_{d})$ can be viewed as an algebroid solution of a complex system of PDEs by differentiating Eq. (2). It is a one-to-one correspondence. This observation motivates us to introduce the notion of algebroid mappings.
Let $M,N$ be complex manifolds of complex dimensions $m,n,$ respectively. Let $\mathscr{M}(M)$ denote the field of all meromorphic functions on $M.$ We give the following definition of algebroid mappings (see Definition 2.4 in Section 2.1).
Definition. We say that a multi-valued mapping $F:M\to N$ is an algebroid mapping, if for every point $y\in f(x)$ with all $x\in M,$ $\xi_{1}\circ F,\cdots,\xi_{n}\circ F$ are local solutions of a system of $n$ independent algebraic equations
$P_{i}\big(x,W_{1},\cdots,W_{n}\big)=0,\ \ \ \ i=1,\cdots,n$
in $W_{1},\cdots,W_{n}$ over $\mathscr{M}(M),$ where $\xi=(\xi_{1},\cdots,\xi_{n})$ is a local holomorphic coordinate near $y.$
According to the definition, we can see that algebroid mappings generalize the notions of meromorphic mappings and algebroid functions.
Examples
  • A 2-valued rational algebroid mapping
    $F=\left[\frac{\zeta_{1}}{\zeta_{0}}:\frac{\zeta_{0}\zeta_{2}}{\zeta_{1}^{2}}:\frac{\sqrt{\zeta_{2}}}{\sqrt{\zeta_{1}}}\right]:\ \mathbb{P}^{2}(\mathbb{C})\to\mathbb{P}^{2}(\mathbb{C}).$
  • A 6-valued transcendental algebroid curve
    $F=\left[z:\sqrt{z}:\sqrt[3]{z}:e^{z}\right]:\ \mathbb{C}\to\mathbb{P}^{3}(\mathbb{C}).$
  • A 12-valued transcendental algebroid mapping
    $F=\big[z_{1}:\sin z_{2}:\sqrt{z_{1}}\sqrt[4]{z_{2}}:\sqrt[3]{z_{3}}e^{z_{4}}\big]:\ \mathbb{C}^{4}\to\mathbb{P}^{3}(\mathbb{C}).$
It is well-known that Nevanlinna theory [17] plays a useful role in the study of meromorphic and algebroid solutions of complex ODEs (see [8, 9, 20]). It seems helpful to develop a value distribution theory of algebroid mappings, as a theoretical tool of investigating algebroid solutions of a complex system of PDEs. Since R. Nevanlinna established Nevanlinna theory, many scholars such as G. Rémoundos, H. Selberg, E. Ullrich and G. Valiron, etc. have been exploring the value distribution theory of algebroid functions. For instance, G. Rémoundos [16] proved a Picard’s theorem which says that a nonconstant $\nu$-valued algebroid function on $\mathbb{C}$ omits $2\nu$ distinct values at most; and then, G. Valiron [29] showed without details a five-value type theorem which says that two nonconstant $\nu$-valued algebroid functions on $\mathbb{C}$ are identical if they share $4\nu$+1 distinct values ignoring multiplicities. For more details, we refer the reader to J. Noguchi [11], Niino-Ozawa [13, 14], J. Suzuki [21], L. Selberg [22, 23], N. Toda [26, 27] and E. Ullrich [28]; and see Hu-Yang [7], He-Xiao [8], M. Ru [18] and Sun-Gao [20], etc.
However, there is very little literature about the value distribution theory of multi-valued analytic mappings, although the researchers have a relatively profound understanding of single-valued analytic mappings (see [12, 17, 19]). In this paper, we shall explore equidistribution theory of algebroid mappings from a complete Kähler manifold into a complex projective manifold. This is a generalization of Carlson-Griffiths equidistribution theory of holomorphic mappings from $\mathbb{C}^{m}$ into a complex projective manifold (see [2]).
1.2. Main Results
Let $M$ be a complex manifold, and $N$ a complex projective manifold.
Theorem 1.1 (=Theorem 3.5).
A $\nu$-valued algebroid mapping $F:M\to N$ defines a $\nu$-leaf complex manifold $\mathcal{M}_{F},$ which is a $\nu$-sheeted ramified analytic covering of $M.$
Corollary 1.2 (=Corollary 3.6).
An algebroid mapping $F:M\to N$ can be lifted to a meromorphic mapping $f:\mathcal{M}_{F}\to N$ via the natural projection $\pi:\mathcal{M}_{F}\to M$ such that $F=f\circ\pi^{-1}.$
Let $(M,\alpha)$ be an $m$-dimensional complete non-compact Kähler manifold with Ricci form $\mathscr{R},$ and $(N,\omega)$ an $n$-dimensional Kähler projective manifold. Let $F:M\to N$ be a $\nu$-valued algebroid mapping. Referring to Sections 6.1 and 6.3, one can define Nevanlinna’s functions $T_{F}(r,\omega),m_{F}(r,D),\overline{N}_{F}(r,D)$ and $T(r,\mathscr{R})$ under different curvature conditions. Since $N$ is projective, one can embed $N$ holomorphically into a complex projective space with Fubini-Study form $\omega_{FS}.$ Let $\omega_{FS}|_{N}$ be the restriction of $\omega_{FS}$ to $N.$ Set
$\left[\frac{\omega_{FS}|_{N}}{\omega}\right]=\inf\big\{s>0:\ \omega_{FS}|_{N}\leq s\omega\big\}.$
$a)$ $M$ has non-positive sectional curvature. Set
(3) $\kappa(r)=\frac{1}{2m-1}\inf_{x\in B(r)}R(x),$
where
$R(x)=\inf_{X\in T_{x}M,\ \|X\|=1}{\rm{Ric}}(X,X)$
is the pointwise lower bound of Ricci curvature of $M,$ and $B(r)$ is the geodesic ball in $M$ centered at some reference point $o$ with radius $r.$
Theorem 1.3 (=Theorem 6.3).
Let $M$ be a complete non-compact Kähler manifold with non-positive sectional curvature. Let $(N,\omega)$ be a Kähler projective manifold of dimension not greater than that of $M.$ Let $D_{1},\cdots,D_{q}$ be effective divisors in general position on $N$ such that each $D_{j}$ is cohomologous to $\omega.$ Let $F:M\rightarrow N$ be a differentiably non-degenerate $\nu$-valued algebroid mapping. Assume that $q\omega-{\rm Ric}(\omega^{n})>0.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$\left(q-(2\nu-2)\left[\frac{\omega_{FS}|_{N}}{\omega}\right]\right)T_{F}(r,\omega)+T_{F}(r,K_{N})+T(r,\mathscr{R})$
$\leq$ $\sum_{j=1}^{q}\overline{N}_{F}(r,D_{j})+O\big(\log T_{F}(r,\omega)+\sqrt{-\kappa(r)}r+\delta\log r\big)$
[…]
Corollary 1.4 (=Corollary 6.4).
Let $M$ be a complete non-compact Kähler manifold with non-positive sectional curvature. Let $H_{1},\cdots,H_{q}$ be hyperplanes of $\mathbb{P}^{n}(\mathbb{C})$ in general position. Let $F:M\rightarrow\mathbb{P}^{n}(\mathbb{C})$ be a differentiably non-degenerate $\nu$-valued algebroid mapping. Assume that $m\geq n.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$\left(q-2\nu-n+1\right)T_{F}(r,\omega_{FS})+T(r,\mathscr{R})$
$\leq$ $\sum_{j=1}^{q}\overline{N}_{F}(r,H_{j})+O\big(\log T_{F}(r,\omega_{FS})+\sqrt{-\kappa(r)}r+\delta\log r\big)$
[…]
$b)$ $M$ is non-parabolic, with non-negative Ricci curvature. Set
(4) $H(r)=\frac{V(r)}{r^{2}}\int_{r}^{\infty}\frac{tdt}{V(t)},$
where $V(r)$ denotes the Riemannian volume of the geodesic ball centered at $o$ with radius $r.$
Theorem 1.5 (=Theorem 6.5).
Let $M$ be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let $(N,\omega)$ be a Kähler projective manifold of dimension not greater than that of $M.$ Let $D_{1},\cdots,D_{q}$ be effective divisors in general position on $N$ such that each $D_{j}$ is cohomologous to $\omega.$ Let $F:M\rightarrow N$ be a differentiably non-degenerate $\nu$-valued algebroid mapping. Assume that $q\omega-{\rm Ric}(\omega^{n})>0.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$\left(q-(2\nu-2)\left[\frac{\omega_{FS}|_{N}}{\omega}\right]\right)T_{F}(r,\omega)+T_{F}(r,K_{N})+T(r,\mathscr{R})$
$\leq$ $\sum_{j=1}^{q}\overline{N}_{F}(r,D_{j})+O\big(\log T_{F}(r,\omega)+\log H(r)+\delta\log r\big)$
[…]
Corollary 1.6 (=Corollary 6.6).
Let $M$ be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let $H_{1},\cdots,H_{q}$ be hyperplanes of $\mathbb{P}^{n}(\mathbb{C})$ in general position. Let $F:M\rightarrow\mathbb{P}^{n}(\mathbb{C})$ be a differentiably non-degenerate $\nu$-valued algebroid mapping. Assume that $m\geq n.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$\left(q-2\nu-n+1\right)T_{F}(r,\omega_{FS})+T(r,\mathscr{R})$
$\leq$ $\sum_{j=1}^{q}\overline{N}_{F}(r,H_{j})+O\big(\log T_{F}(r,\omega_{FS})+\log H(r)+\delta\log r\big)$
[…]
2. Algebroid Mappings
2.1. Algebroid Mappings
Let $M$ be a complex manifold of complex dimension $m.$ Let us recall that a $\nu$-valued algebroid function $W$ on $M$ is defined by an irreducible algebraic equation of degree $\nu$
$A_{\nu}(x)W^{\nu}+A_{\nu-1}(x)W^{\nu-1}+\cdots+A_{0}(x)=0,\ \ \ \ A_{\nu}\not\equiv 0$
with holomorphic functions $A_{0},\cdots,A_{\nu}$ locally defined on $M,$ such that they well define a meromorphic mapping $A_{0}:\cdots:A_{\nu}]:M\rightarrow\mathbb{P}^{\nu}(\mathbb{C})$; or in other words, it is defined by an irreducible algebraic equation of degree $\nu$
$W^{\nu}+B_{\nu-1}(x)W^{\nu-1}+\cdots+B_{0}(x)=0$
over the field $\mathscr{M}(M)$ of all meromorphic functions on $M.$ Note that
$B_{j}=\frac{A_{j}}{A_{\nu}},\ \ \ \ j=0,\cdots,\nu-1.$
It is well-known that algebroid functions extend the notions of meromorphic functions and algebraic functions.
Let us give any two polynomials in $W$:
$P_{1}(x,W)$ $=$ $W^{\nu_{1}}+B_{1(\nu_{1}-1)}(x)W^{\nu_{1}-1}+\cdots+B_{10}(x),$
$P_{2}(x,W)$ $=$ $W^{\nu_{2}}+B_{2(\nu_{2}-1)}(x)W^{\nu_{2}-1}+\cdots+B_{20}(x)$
over $\mathscr{M}(M).$ The resultant of $P_{1}(x,W)$ and $P_{2}(x,W)$ is defined by
$R=\left|\begin{matrix}1&B_{1(\nu_{1}-1)}&\cdots&B_{12}&B_{11}&B_{10}&0&0&\cdots&0\\ 0&1&\cdots&B_{13}&B_{12}&B_{11}&B_{10}&0&\cdots&0\\ \vdots&\vdots&&\vdots&\vdots&\vdots&\vdots&\vdots&&\vdots\\ 0&0&\cdots&1&B_{1(\nu_{1}-1)}&B_{1(\nu_{1}-2)}&B_{1(\nu_{1}-3)}&B_{1(\nu_{1}-4)}&\cdots&B_{10}\\ 1&B_{2(\nu_{2}-1)}&\cdots&B_{22}&B_{21}&0&0&0&\cdots&0\\ 0&1&\cdots&B_{23}&B_{22}&B_{21}&0&0&\cdots&0\\ \vdots&\vdots&&\vdots&\vdots&\vdots&\vdots&\vdots&&\vdots\\ 0&0&\cdots&0&1&B_{2(\nu_{2}-1)}&B_{2(\nu_{2}-2)}&B_{2(\nu_{2}-3)}&\cdots&B_{21}\\ \end{matrix}\right|.$
It is known that $R$ is obtained based on Sylvester’s elimination method. Let $\alpha_{1}(x),\cdots,\alpha_{\nu_{1}}(x)$ and $\beta_{1}(x),\cdots,\beta_{\nu_{2}}(x)$ be two groups of roots of $P_{1}(x,W)$ and $P_{2}(x,W),$ respectively. Then
$R=\prod_{1\leq i\leq\nu_{1},1\leq j\leq\nu_{2}}(\alpha_{i}-\beta_{j}).$
Denote by $\mathscr{A}(M)$ the set of all algebroid functions on $M.$ We have:
Theorem 2.1.
Under the usual multiplication and addition, $\mathscr{A}(M)$ forms a field with identity element $1$ and zero element $0.$
Proof.
We first show that $\mathscr{A}(M)$ is closed under the usual multiplication and addition. Let $W_{1},W_{2}\in\mathscr{A}(M)$ be defined by irreducible algebraic equations $P_{1}(x,W_{1}),P_{2}(x,W_{2}),$ respectively. Set $W=W_{1}W_{2}.$ Substitute $W_{2}=W/W_{1}$ into $P_{2}(x,W_{2}),$ we obtain $P_{2}(x,W/W_{1})=0.$ Regard
$Q(x,W_{1}):=W_{1}^{\nu_{2}}P_{2}(x,W/W_{1})$
as a polynomial in $W_{1},$ where $\nu_{2}=\deg P_{2}.$ Applying Sylvester’s elimination method, we deduce that $R(x,W)=0,$ where $R$ is the resultant of $P_{1}(x,W_{1})$ and $Q(x,W_{1}).$ It means that $\mathscr{A}(M)$ is closed under the usual multiplication. Similarly, $\mathscr{A}(M)$ is closed under the usual addition. Secondly, we prove that $1,0$ are identity element and zero element of $\mathscr{A}(M),$ respectively. It is clear that $1,0\in\mathscr{A}(M).$ For any $W\in\mathscr{A}(M)$ with defining equation
$P(x,W):=W^{\nu}+B_{\nu-1}W^{\nu-1}+\cdots+B_{0}=0$
over $\mathscr{M}(M),$ it suffices to confirm that $W\in\mathscr{A}(M)$ and $1/W\in\mathscr{A}(M)$ for $W\not\equiv 0.$ We see that W-W satisfies the algebraic equation
$(-W)^{\nu}+\tilde{B}_{\nu-1}(-W)^{\nu-1}+\cdots+\tilde{B}_{0}=0,$
where $\tilde{B}_{j}=(-1)^{\nu-j}B_{j}$ with $j=0,\cdots,\nu-1.$ Hence, we have $W_{1}\in\mathscr{A}(M).$ Also, we have $1/W\in\mathscr{A}(M)$ for $W\not\equiv 0,$ due to $W^{-\nu}P(x,1/W)=0.$ Finally, the commutative law, associative law and distributive law can be examined also easily. This completes the proof. ∎
Corollary 2.2.
$\mathscr{A}(M)$ is an algebraically closed field, and further an algebraic closure of $\mathscr{M}(M).$
Proof.
At first, we prove that $\mathscr{A}(M)$ is an algebraically closed field, i.e., any $W$ defined by an irreducible algebraic equation
$C_{\nu}(x)W^{\nu}+C_{\nu-1}(x)W^{\nu-1}+\cdots+C_{0}(x)=0,\ \ \ \ C_{\nu}\not\equiv 0$
over $\mathscr{A}(M)$ is an element of $\mathscr{A}(M).$ Using Theorem 2.1, the above equation can be written as
(5) $W^{\nu}+\tilde{C}_{\nu-1}(x)W^{\nu-1}+\cdots+\tilde{C}_{0}(x)=0,$
[…]
Let $N$ be a complex manifold of complex dimension $n.$ We shall introduce the notion of algebroid mappings as follows.
Definition 2.3.
A multi-valued mapping $F:M\to N$ is called an algebroid mapping, if for every point $y\in f(x)$ with all $x\in M,$ $\xi_{1}\circ F,\cdots,\xi_{n}\circ F$ are local solutions of a system of $n$ independent algebraic equations
$P_{i}\big(x,W_{1},\cdots,W_{n}\big)=0,\ \ \ \ i=1,\cdots,n$
in $W_{1},\cdots,W_{n}$ over $\mathscr{M}(M),$ where $\xi=(\xi_{1},\cdots,\xi_{n})$ is a local holomorphic coordinate around $y.$
In particular, we treat the case when $N$ is a complex projective manifold. Let $\jmath:N\hookrightarrow\mathbb{P}^{d}(\mathbb{C})$ be the holomorphic embedding. Let $\zeta=[\zeta_{0}:\cdots:\zeta_{d}]$ be the homogenous coordinate of $\mathbb{P}^{d}(\mathbb{C}).$ Then, we follow an extrinsic definition of algebroid mappings:
Definition 2.4.
Assume that $N$ is a complex projective manifold. A multi-valued mapping $F:M\to N$ is called an algebroid mapping, if
$\jmath\circ F=\big[\zeta_{0}\circ\jmath\circ F:\cdots:\zeta_{d}\circ\jmath\circ F\big]:\ M\to\mathbb{P}^{d}(\mathbb{C})$
defines an algebroid mapping, i.e.,
$W_{i}:=\frac{\zeta_{i}\circ\jmath\circ F}{\zeta_{0}\circ\jmath\circ F},\ \ \ \ i=1,\cdots,d$
satisfy a system of $d$ independent algebraic equations
$P_{i}\big(x,W_{1},\cdots,W_{n}\big)=0,\ \ \ \ i=1,\cdots,n$
in $W_{1},\cdots,W_{n}$ over $\mathscr{M}(M).$
According to Corollary 2.2, if $W_{1},\cdots,W_{d}$ are defined by some system of $d$ independent algebraic equations over $\mathscr{M}(M),$ then each $W_{i}$ is an algebroid function on $M.$ Thus, we obtain an alternative, but better extrinsic definition of algebroid mappings as follows.
Definition 2.5.
A $\nu$-valued mapping $F:M\to N$ is called a $\nu$-valued algebroid mapping, if
$\jmath\circ F=\big[\zeta_{0}\circ\jmath\circ F:\cdots:\zeta_{d}\circ\jmath\circ F\big]:\ M\to\mathbb{P}^{d}(\mathbb{C})$
defines a $\nu$-valued algebroid mapping, i.e.,
$W_{i}:=\frac{\zeta_{i}\circ\jmath\circ F}{\zeta_{0}\circ\jmath\circ F},\ \ \ \ i=1,\cdots,d$
are $\nu_{i}$-valued algebroid functions on $M,$ respectively, such that $\nu=\nu_{1}\cdots\nu_{d}.$ Here, $W_{1},\cdots,W_{d}$ are understood to be independent in choosing single-valued components since they are in different coordinate positions.
Combining Definition 2.5 with Theorem 2.1, it is immediate that
Corollary 2.6.
Let $W_{0},\cdots,W_{n}$ be algebroid functions without poles and common zero divisors on $M.$ Then
$F=\big[W_{0}:\cdots:W_{n}\big]:\ M\to\mathbb{P}^{n}(\mathbb{C})$
is an algebroid mapping.
2.2. Quasi-regular Mapping Elements
From now on, we assume always that $N$ is a complex projective manifold. For technical reasons, we equip $M$ with a Hermitian metric. Let $F:M\to N$ be a $\nu$-valued algebroid mapping, i.e., $W_{i}=\zeta_{i}\circ\jmath\circ F/\zeta_{0}\circ\jmath\circ F$ is a $\nu_{i}$-valued algebroid function on $M$ with defining equation
(6) $A_{i\nu_{i}}(x)W^{\nu_{i}}_{i}+A_{i(\nu_{i}-1)}(x)W^{\nu_{i}-1}_{i}+\cdots+A_{i0}(x)=0,\ \ \ \ A_{i\nu_{i}}\not\equiv 0$
for $i=1,\cdots,d.$ Note that $\nu=\nu_{1}\cdots\nu_{d}.$
Definition 2.7.
[…]
For $i=1,\cdots,d,$ put
(7) $\Psi_{i}(x,W_{i}):=A_{i\nu_{i}}(x)W^{\nu_{i}}_{i}+A_{i(\nu_{i}-1)}(x)W^{\nu_{i}-1}_{i}+\cdots+A_{i0}(x).$
Then
$\Psi_{i,i}(x,W_{i})$ $:=$ $\frac{\partial\Psi_{i}(x,W_{i})}{\partial W_{i}}$
$=$ $\nu_{i}A_{i\nu_{i}}(x)W^{\nu_{i}-1}_{i}+(\nu_{i}-1)A_{i(\nu_{i}-1)}(x)W^{\nu_{i}-2}_{i}+\cdots+A_{i1}(x).$
Theorem 2.8.
Let $(x_{0},y_{0})\in M\times N$ be a pair such that
$\Psi_{i}(x_{0},c_{i})=0,\ \ \ \ \Psi_{i,i}(x_{0},c_{i})\not=0,\ \ \ \ i=1,\cdots,d,$
where
$c_{i}=\frac{\zeta_{i}\circ\jmath(y_{0})}{\zeta_{0}\circ\jmath(y_{0})},\ \ \ \ i=1,\cdots,d.$
Then, there exists a unique regular mapping element $(x_{0},H)$ of $F$ such that $y_{0}=H(x_{0}).$ Moreover, $\zeta_{0}\circ\jmath\circ H$ has no zeros.
Proof.
Write
$\Psi_{i}(x,W_{i})$
$=$ $B_{i\nu_{i}}(x)(W_{i}-c_{i})^{\nu_{i}}+B_{i(\nu_{i}-1)}(x)(W_{i}-c_{i})^{\nu_{i}-1}+\cdots+B_{i0}(x)$
$=$ $(W_{i}-c_{i})P_{i}(x,W_{i})+B_{i0}(x),$
where
$P_{i}(x,W_{i})$
$=$ $B_{i\nu_{i}}(x)(W_{i}-c_{i})^{\nu_{i}-1}+B_{i(\nu_{i}-1)}(x)(W_{i}-c_{i})^{\nu_{i}-2}+\cdots+B_{i1}(x).$
By conditions, we have
$B_{i0}(x_{0})$ $=$ $\Psi_{i}(x_{0},c_{i})=0,$
$P_{i}(x_{0},c_{i})$ $=$ $\Psi_{i,i}(x_{0},c_{i}):=b_{i}\not=0.$
Using the continuation, there exist $r,\rho>0$ such that
$|B_{i0}(x)|<\frac{\rho|b_{i}|}{3},\ \ \ \ |P_{i}(x,W_{i})|>\frac{2|b_{i}|}{3},\ \ \ \ i=1,\cdots,d$
for all $x,W_{i}$ satisfying ${\rm{dist}}(x_{0},x)\leq r$ and $|W_{i}-c_{i}|\leq\rho.$ In further, restrict $W_{i}$ to $|W_{i}-c_{i}|=\rho,$ we have for $i=1,\cdots,d$
$\left|(W_{i}-c_{i})P_{i}(x,W_{i})\right|>\frac{2\rho|b_{i}|}{3}>\frac{\rho|b_{i}|}{3}>|B_{i0}(x)|.$
According to Rouché’s theorem, for every $x$ with ${\rm{dist}}(x_{0},x)\leq r,$ we deduce that $\Psi_{i}(x,W_{i})=(W_{i}-c_{i})P_{i}(x,W_{i})+B_{i0}$ as well as that $(W_{i}-c_{i})P_{i}(x,W_{i})$ have the same number of zeros in the disc $\{|W_{i}-c_{i}|<\rho\}.$ In fact, the number of zeros is 1, since $(W_{i}-c_{i})P_{i}(x,W_{i})$ in $W_{i}$ has only one zero $W_{i}=c_{i}$ in the disc $\{|W_{i}-c_{i}|<\rho\}$ with ${\rm{dist}}(x_{0},x)\leq r.$ This implies that for $i=1,\cdots,d,$ $\Psi_{i}(x,W_{i})=0$ uniquely determines a solution $W_{i}=u_{i}(x)$ in a small geodesic ball neighborhood $B(x_{0},r_{0})$ of $x_{0}$ with radius $r_{0}$ such that
$\Psi_{i}(x,u_{i})=0,\ \ \ \ c_{i}=u_{i}(x_{0}).$
Clearly, every $u_{i}$ is continuous on $B(x_{0},r_{0}).$ Again, since $\jmath$ is a holomorphic embedding, it is easy to verify that there exists a unique continuous mapping $H:B(x_{0},r_{0})\to N$ such that $\jmath\circ H=[1:u_{1}:\cdots:u_{d}]$ with $y_{0}=H(x_{0})$ and $u_{i}=\zeta_{i}\circ\jmath\circ H/\zeta_{0}\circ\jmath\circ H$ for $i=1,\cdots,d.$ Since $u_{1},\cdots,u_{d}$ are continuous, it implies that $\zeta_{0}\circ\jmath\circ H$ has no zeros.
Next, we show that $H$ is holomorphic on $B(x_{0},r_{0}).$ One just needs to show without loss of generality that $H$ is analytic at $x_{0}.$ Take a local holomorphic coordinate $z=(z_{1},\cdots,z_{m})$ near $x_{0}$ such that $z(x_{0})=\textbf{0}.$ Thanks to Hartog’s theorem, it is sufficient to show that $u_{i}$ is holomorphic in every holomorphic direction $z_{k}$ at 0 for $i=1,\cdots,d.$ Put
$u_{i,k}(z_{k})$ $=$ $u_{i}(0,\cdots,z_{k},\cdots,0),$
$B_{ij,k}(z_{k})$ $=$ $B_{ij}(0,\cdots,0,z_{k},0,\cdots,0),$
where $i=1,\cdots,d,$ $j=0,\cdots,\nu_{i},$ and $k=1,\cdots,m.$ Note that
$\frac{u_{i,k}-c_{i}}{z_{k}}$
$=$ $-\frac{B_{i0,k}}{z_{k}}\frac{1}{B_{i\nu_{i},k}(u_{i,k}-c_{j})^{\nu_{i}-1}+B_{i(\nu_{i}-1),k}(u_{i,k}-c_{i})^{\nu_{i}2}+\cdots+B_{i1,k}}$
$\rightarrow$ $-\frac{1}{B_{i1}(\textbf{0})}\frac{\partial B_{i0}}{\partial z_{k}}(\textbf{0})$
as $z_{k}\rightarrow 0.$ That is,
$\frac{\partial u_{i,k}}{\partial z_{k}}(\textbf{0})=-\frac{1}{B_{i1}(\textbf{0})}\frac{\partial B_{i0}}{\partial z_{k}}(\textbf{0}).$
Thus, $u_{i}$ is holomorphic in the holomorphic direction $z_{k}$ at $\textbf{0}.$ This completes the proof. ∎
Definition 2.9.
Let $x_{0}$ be an arbitrary point in $M.$ We call $x_{0}$ a critical point of $W_{i},$ if either $A_{i\nu_{i}}(x_{0})=0$ or $\Psi_{i}(x_{0},W_{i})$ has a multiple root as a polynomial in $W_{i}.$ We use $\mathscr{S}_{W_{i}}$ to denote the set of all critical points of $W_{i},$ called the critical set of $W_{i}.$ A regular point of $W_{i}$ means a point in $\mathscr{T}_{W_{i}}:=M\setminus\mathscr{S}_{W_{i}},$ where $\mathscr{T}_{W_{i}}$ is called the regular set of $W_{i}.$ Denote by $\mathscr{M}_{W_{i}}$ the set of all points $x_{0}$ in $\mathscr{S}_{W_{i}}$ such that $\Psi_{i}(x_{0},W_{i})$ has a multiple root as a polynomial in $W_{i},$ where $\mathscr{M}_{W_{i}}$ is called the multiple set of $W_{i}$ and a point in $\mathscr{M}_{W_{i}}$ is called a multiple point of $W_{i}.$
For $x_{0}\in\mathscr{T}_{M_{i}},$ the fundamental theorem of algebra says that the algebraic equation
$A_{i\nu_{i}}(x_{0})W^{\nu_{i}}_{i}+A_{i(\nu_{i}-1)}(x_{0})W_{i}^{\nu_{i}-1}+\cdots+A_{i0}(x_{0})=0$
has exactly $\nu_{i}$ distinct complex roots. Hence, Theorem 2.8 yields that
Corollary 2.10.
For each $x\in M\setminus\cup_{i=1}^{d}\mathscr{S}_{W_{i}},$ there exist exactly $\nu$ distinct regular mapping elements $(x,H_{1}),\cdots,(x,H_{\nu})$ of $F.$ Moreover, each $\zeta_{0}\circ\jmath\circ H_{j}$ has no zeros.
To find the common roots of $\Psi_{i}(x,W_{i}),\Psi_{i,i}(x,W_{i}),$ we treat the resultant $R_{\Psi_{i}}$ of $\Psi_{i}(x,W_{i}),\Psi_{i,i}(x,W_{i}),$ i.e.,
$R_{\Psi_{i}}=\left|\begin{matrix}A_{i\nu_{i}}&A_{i(\nu_{i}-1)}&\cdots&A_{i2}&A_{i1}&A_{i0}&0&0&\cdots&0\\ 0&A_{i\nu_{i}}&\cdots&A_{i3}&A_{i2}&A_{i1}&A_{i0}&0&\cdots&0\\ \vdots&\vdots&&\vdots&\vdots&\vdots&\vdots&\vdots&&\vdots\\ 0&0&\cdots&A_{i\nu_{i}}&A_{i(\nu_{i}-1)}&A_{i(\nu_{i}-2)}&A_{i(\nu_{i}-3)}&A_{i(\nu_{i}-4)}&\cdots&A_{i0}\\ B_{i\nu_{i}}&B_{i(\nu_{i}-1)}&\cdots&B_{i2}&B_{i1}&0&0&0&\cdots&0\\ 0&B_{i\nu_{i}}&\cdots&B_{i3}&B_{i2}&B_{i1}&0&0&\cdots&0\\ \vdots&\vdots&&\vdots&\vdots&\vdots&\vdots&\vdots&&\vdots\\ 0&0&\cdots&0&B_{i\nu_{i}}&B_{i(\nu_{i}-1)}&B_{i(\nu_{i}-2)}&B_{i(\nu_{i}-3)}&\cdots&B_{i1}\\ \end{matrix}\right|,$
where $B_{ij}=iA_{ij}$ for $i=1,\cdots,\nu_{i}.$ Applying Sylvester’s elimination method, we see that $R_{\Psi_{i}}(x)=0$ if and only if $A_{i\nu_{i}}(x)=0$ or $\Psi_{i}(x,W_{i})$ has a multiple root as a polynomial in $W_{i},$ i.e., $x\in\mathscr{S}_{W_{i}}.$ Therefore, $\mathscr{S}_{W_{i}}$ is an analytic set of complex codimension $1.$ Define the discriminant of $\Psi_{i}(x,W_{i})$ by
$J_{\Psi_{i}}(x)=(-1)^{\frac{\nu_{i}(\nu_{i}-1)}{2}}\frac{R_{\Psi_{i}}(x)}{A_{i\nu_{i}}(x)}\not\equiv 0,\ \ \ \ i=1,\cdots,d.$
Write
$\Psi_{i}(x,W_{i})=A_{i\nu_{i}}(x)(W_{i}-w_{i1}(x))\cdots(W_{i}-w_{i\nu_{i}}(x)).$
A direct computation leads to
$J_{\Psi_{i}}(x)=A_{i\nu_{i}}(x)^{2\nu_{i}-2}\prod_{1\leq j<k\leq\nu_{i}}\big(w_{ij}(x)-w_{ik}(x)\big)^{2}.$
If $x_{0}\in\mathscr{S}_{W_{i}}\setminus A^{-1}_{i\nu_{i}}(0),$ then $\Psi_{i}(x_{0},W_{i})$ has a multiple root as a polynomial in $W_{i}.$ If $x_{0}\in A^{-1}_{i\nu_{i}}(0)\setminus I_{\mathscr{A}_{F}},$ then we set $V_{i}=1/W_{i}$ and rewrite (6) as
$\Psi_{i}(x,W_{i})=V^{-\nu_{i}}_{i}\Phi_{i}(x,V_{i}),$
where
$\Phi_{i}(x,V_{i})=A_{i\nu_{i}}(x)+A_{i(\nu_{i}-1)}(x)V_{i}+\cdots+A_{i0}(x)V^{\nu_{i}}_{i}.$
Since $\Phi_{i}(x,0)=A_{i\nu_{i}}(x),$ we can deduce that $x_{0}$ is a zero of $V_{i},$ i.e., a pole of $W_{i}.$ In fact, it yields from Theorem 4.1 (see Section 2.3) that $J_{\Psi_{i}}(x_{0})=0$ if and only if $x_{0}\in\mathscr{M}_{W_{i}}.$ Thus, Corollary 2.10 can extend to $M\setminus\cup_{i=1}^{d}\mathscr{M}_{W_{i}}.$
Corollary 2.11.
For each $x\in M\setminus\cup_{i=1}^{d}\mathscr{M}_{W_{i}},$ there exist exactly $\nu$ distinct quasi-regular mapping elements $(x,H_{1}),\cdots,(x,H_{\nu})$ of $F.$
Let $F_{1},\cdots,F_{\nu}$ stand for the $\nu$ distinct single-valued components of $F.$ We call $x_{0}\in M$ a multiple point of $F,$ if $F_{i}(x_{0})=F_{j}(x_{0})$ for some $i\not=j.$ Denote by $\mathscr{M}_{F}$ the set of all multiple points of $F,$ which is called the multiple set of $F.$
Lemma 2.12.
We have
$\mathscr{M}_{F}=\cup_{i=1}^{d}\mathscr{M}_{W_{i}}.$
Proof.
If $x_{0}\in\mathscr{M}_{W_{k}}$ for some $k$ with $1\leq k\leq d,$ then there exist two distinct single-valued components $W_{ki},W_{kj}$ of $W_{k}$ such that $W_{ki}(x_{0})=W_{kj}(x_{0}).$ We may assume without loss of generality that $k=1.$ Let $W_{l1}$ be a single-valued component of $W_{l}$ for $l=2,\cdots,\nu.$ Take two single-valued components $F_{i},F_{j}$ of $F$ such that $\jmath\circ F_{p}=[1:W_{1p}:W_{21}:\cdots:W_{\nu 1}]$ with $p=i,j.$ Clearly, we have $\jmath\circ F_{i}(x_{0})=\jmath\circ F_{j}(x_{0}).$ Since $\jmath$ is injective, we obtain $F_{i}(x_{0})=F_{j}(x_{0}),$ which implies that $x_{0}\in\mathscr{M}_{F}.$ If $x_{0}\in\mathscr{M}_{F},$ then we have $F_{i}(x_{0})=F_{j}(x_{0})$ for some two distinct single-valued components $F_{i},F_{j}$ of $F.$ It yields that
$\frac{\zeta_{k}\circ\jmath\circ F_{i}}{\zeta_{0}\circ\jmath\circ F_{i}}=\frac{\zeta_{k}\circ\jmath\circ F_{j}}{\zeta_{0}\circ\jmath\circ F_{j}},\ \ \ \ k=1,\cdots,d.$
This gives that $x_{0}\in\cup_{i=1}^{d}\mathscr{M}_{W_{i}}.$ The proof is completed. ∎
By Corollary 2.11 and Lemma 2.12, we have:
Corollary 2.13.
For each $x\in M\setminus\mathscr{M}_{F},$ there exist exactly $\nu$ distinct quasi-regular mapping elements $(x,H_{1}),\cdots,(x,H_{\nu})$ of $F.$
2.3. Analytic Continuation of Quasi-regular Mapping Elements
We review that a quasi-regular mapping element of $F$ means a pair $(x,H)$ such that $\jmath\circ H$ is a meromorphic mapping in a neighborhood $U(x)$ of $x.$ For convenience, we write
$\tilde{x}=(x,H)=(U(x),H)$
without any confusion. Instead of $U(x),$ a geodesic ball neighborhood $B(x,r)$ centered at $x$ with radius $r$ is usually considered.
Let $\tilde{x}_{1}=(U(x_{1}),H_{1}),$ $\tilde{x}_{2}=(U(x_{2}),H_{2})$ be any two quasi-regular mapping elements of $F.$ Define an equivalent relation as follows: $\tilde{x}_{1}\sim\tilde{x}_{2}$ if and only if
$x_{1}=x_{2};\ \ \ \ H_{1}(x)=H_{2}(x),\ \ \ \ ^{\forall}x\in U(x_{1})\cap U(x_{2}).$
Set
$\mathcal{M}^{0}_{F}=\big\{\text{all quasi-regular mapping elements of $F$}\big\}/\sim.$
A quasi-regular mapping element $\tilde{y}=(y,G)$ is said to be a direct analytic continuation of a quasi-regular mapping element $\tilde{x}=(U(x),H),$ if $y\in U(x)$ and $G=H$ in a neighborhood of $y.$ Using the analytic continuation, we can regard that
$\tilde{y}=(y,G)=(y,H).$
A neighborhood V(x~)V($\tilde${x}) of $\tilde{x}$ is defined by a set (containing $\tilde{x}$) of all $\tilde{y}\in\mathcal{M}^{0}$ such that $\tilde{y}$ is a direct analytic continuation of $\tilde{x}$ satisfying $y\in U_{1}(x),$ where $U_{1}(x)$ is a neighborhood of $x$ in $U(x).$ Namely, $V(\tilde{x})=\{(y,H):y\in U_{1}(x)\}.$ In particular, a $r$-neighborhood $V(\tilde{x},r)$ of $\tilde{x}$ (satisfying that $B(x,r)\subset U(x)$), is defined by the set (which contains $\tilde{x}$) of all $\tilde{y}\in\mathcal{M}^{0}$ such that $\tilde{y}$ is a direct analytic continuation of $\tilde{x}$ satisfying that $y\in B(x,r).$ In other words, it is defined to be $V(\tilde{x},r)=\{(y,H):y\in B(x,r)\}.$ A subset $\mathcal{E}\subseteq\mathcal{M}^{0}$ is called an open set, if either $\mathcal{E}$ is an empty set or every point $\tilde{x}\in\mathcal{E}$ is an inner point, i.e., there exists a $\epsilon$-neighborhood $V(\tilde{x},\epsilon)$ of $\tilde{x}$ such that $V(\tilde{x},\epsilon)\subset\mathcal{E}.$ It is not hard to check that it defines a topology of $\mathcal{M}^{0}.$
Theorem 2.14.
$\mathcal{M}^{0}_{F}$ is a Hausdorff space.
Proof.
Pick $\tilde{x}_{1}\not=\tilde{x}_{2}\in\mathcal{M}^{0}_{F}.$ If $x_{1}\not=x_{2},$ then we can consider $\epsilon$-neighborhoods $V(\tilde{x}_{1},\epsilon)$ of $\tilde{x}_{1}$ and $V(\tilde{x}_{2},\epsilon)$ of $\tilde{x}_{2}$ with $0<\epsilon<{\rm{dist}}(x_{1},x_{2})/2.$ It is clear that $V(\tilde{x}_{1},\epsilon)\cap V(\tilde{x}_{2},\epsilon)=\emptyset.$ If $V(\tilde{x}_{1},r_{1})\cap V(\tilde{x}_{2},r_{2})\not$$\emptyset$ for $x_{1}=x_{2},$ then we can take a quasi-regular mapping element $(x_{0},H_{0})\in V(\tilde{x}_{1},\epsilon)\cap V(\tilde{x}_{2},\epsilon),$ which gives $H_{1}=H_{0}=H_{2}$ in some smaller neighborhood of $x_{0}.$ By the uniqueness theorem of analytic mappings, we obtain $H_{1}\equiv H_{2}.$ However, it contradicts with $\tilde{x}_{1}\not=\tilde{x}_{2}.$
Definition 2.15.
Let $\tilde{x}$, $\tilde{y}$ be two quasi-regular mapping elements of $F.$ Let $\gamma:[0,1]\to M$ be a curve with $x=\gamma(0)$ and $y=\gamma(1).$ We say that $\tilde{y}$ is an analytic continuation of $\tilde{x}$ along $\gamma,$ if
$(i)$ for any $t\in[0,1],$ there is a quasi-regular mapping element $\tilde{\gamma}(t)$ of $F;$
$(ii)$ for any $t_{0}\in[0,1]$ and any $\epsilon>0,$ there exists $\delta>0$ such that $|\gamma(t)-\gamma(t_{0})|<\epsilon$ for all $t\in[0,1]\cap\{|t-t_{0}|<\delta\},$ and that $\tilde{\gamma}(t)$ is a direct analytic continuation of $\tilde{\gamma}(t_{0}).$
Remark 2.16.
An analytic continuation of elements in $\mathcal{M}^{0}_{F}$ along a curve $\gamma$ in $M$ is a continuous mapping of $\gamma$ into $\mathcal{M}^{0}_{F},$ i.e., an analytic continuation of $\tilde{x}$ to $\tilde{y}$ along $\gamma$ is a curve connecting $x$ and $y.$ Moreover, if $\tilde{y}$ is an analytic continuation of $\tilde{x}$ along a curve, then $\tilde{x}$ is also an analytic continuation of $\tilde{y}$ along the same curve. Thus, the analytic continuation is symmetric.
Theorem 2.17 (Analytic continuation).
Let $\gamma:[0,1]\to M\setminus\mathscr{M}_{F}$ be a curve. Then, any mapping element $\tilde{\gamma}(0)\in\mathcal{M}^{0}_{F}$ can continue analytically to $\gamma(1)$ along $\gamma$ and attain to a unique mapping element $\tilde{\gamma}(1)\in\mathcal{M}^{0}_{F}.$
Proof.
Set
$\tau=\sup\big\{t\in[0,1]:\ \text{$\tilde{\gamma}(0)$ can continue analytically to a $(\gamma(t),H_{t})\in\mathcal{M}^{0}_{F}$}\big\}.$
It is clear that $0<\tau\leq 1.$ It is sufficient to confirm that $\tau=1.$ Otherwise, we may assume that $\tau<1.$ Using Corollary 2.13, there exist exactly $\nu$ mapping elements of $F$ at $\gamma(\tau),$ saying
$\left(B(\gamma(\tau),r_{\tau}),G_{1}\right),\cdots,\left(B(\gamma(\tau),r_{\tau}),G_{\nu}\right)\in\mathcal{M}^{0}_{F}.$
Take $\delta>0$ small enough such that $\{\gamma(t):t\in[\tau-\delta,\tau+\delta]\}\subset B(\gamma(\tau),r_{\tau})$ with $\delta<\mathop{\mathrm{min}}\{\tau,1-\tau\}.$ Then, there is a unique integer $k_{0}\in\{1,\cdots,\nu\}$ such that $H_{\tau-\delta}=G_{k_{0}}$ in a neighborhood of $\gamma(\tau-\delta).$ Notice that $G_{k_{0}}$ is defined on $B(\gamma(\tau),r_{\tau}),$ then for any $t\in[\tau-\delta,\tau+\delta],$ one can redefine $(\gamma(t),G_{k_{0}})\in\mathcal{M}^{0}_{F}.$ Combined with the definition of $\{\gamma(t),H_{t}\}_{0\leq t\leq\tau-\delta},$ it gives immediately that $\tilde{\gamma}(0)$ can continue analytically to $\gamma(\tau+\delta)$ along $\gamma,$ i.e., $(\gamma(\tau+\delta),G_{k_{0}})\in\mathcal{M}^{0}_{F}$ is an analytic continuation of $\tilde{\gamma}(0)$ along $\gamma.$ However, it contradicts with the definition of $\tau.$ Hence, we obtain $\tau=1.$ The uniqueness of $\tilde{\gamma}(1)$ is immediate due to the uniqueness theorem of analytic mappings. ∎
Corollary 2.18.
Let $\tilde{X}=\{(x,H_{j})\}_{j=1}^{\nu},\tilde{Y}=\{(y,G_{j})\}_{j=1}^{\nu}\subset\mathcal{M}^{0}_{F}$ be two groups of distinct mapping elements at $x,y\in M\setminus\mathscr{M}_{F},$ respectively. Then, $\tilde{X},\tilde{Y}$ can continue analytically to each other along any curve in $M\setminus\mathscr{M}_{F}.$
Theorem 2.19 (Connectivity).
Any two quasi-regular mapping elements in $\mathcal{M}^{0}_{F}$ can continue analytically to each other along a curve in $M\setminus\mathscr{M}_{F}.$
Proof.
Let $\{(x,H_{j})\}_{j=1}^{\nu}\subset\mathcal{M}^{0}_{F}$ be $\nu$ distinct quasi-regular mapping elements at any point $x\in M\setminus\mathscr{M}_{F}.$ Using Corollary 2.18, it is sufficient to prove that $(x,H_{1}),\cdots,(x,H_{\nu})$ can continue analytically to each other. Let $n$ stand for the largest integer such that there exist $n$ mapping elements in $\{(x,H_{j})\}_{j=1}^{\nu},$ which can continue analytically to each other along a curve. Clearly, we have $1\leq n\leq\nu.$ Without loss of generality, we shall assume that these $n$ mapping elements are $(x,H_{1}),\cdots,(x,H_{n}).$ For $j=1,\cdots,n,$ set
$u_{ji}=\zeta_{i}\circ\jmath\circ H_{j},\ \ \ \ i=1,\cdots,d.$
From the assumption, $n$ is also the largest integer such that the meromorphic mappings $1:u_{j1}:\cdots:u_{jd}]$ into $\mathbb{P}^{d}(\mathbb{C})$ defined in a neighborhood of $x,$ with $j=1,\cdots,n,$ can continue analytically to each other along a curve. Write
$(W-u_{1i})(W-u_{2i})\cdots(W-u_{ni})=W^{n}+B_{(n-1)i}W^{n-1}+\cdots+B_{0i},$
where $B_{0i},\cdots,B_{(n-1)i}$ are holomorphic for $i=1,\cdots,d.$ By Vieta’s theorem, we obtain
$\sum_{1\leq j\leq n}u_{ji}$ $=$ $-B_{(n-1)i}$
$\cdots\cdots\cdots$
$\sum_{1\leq j_{1}<\cdots<j_{k}\leq n}u_{j_{1}i}\cdots u_{j_{k}i}$ $=$ $(-1)^{k}B_{(n-k)i}$
$\cdots\cdots\cdots$
$u_{1i}\cdots u_{ni}$ $=$ $(-1)^{n}B_{0i}$
It suffices to show that
$(i)$ $B_{1i},\cdots,B_{ni}$ can continue analytically to the whole $M\setminus\mathscr{M}_{F};$
$(ii)$ $n=\nu.$
Using Theorem 2.17, $\{(x,H_{j})\}_{j=1}^{\nu}$ can continue analytically to $\{(y,G_{j})\}_{j=1}^{\nu}$ along a curve for each $y\in M\setminus\mathscr{M}_{F},$ where $(G_{1},\cdots,G_{n})$ is a permutation of $(H_{1},\cdots,H_{n})$ due to the assumption. For $j=1,\cdots,n,$ set
$v_{ji}=\zeta_{i}\circ\jmath\circ G_{j},\ \ \ \ i=1,\cdots,d.$
Then, $(v_{1i},\cdots,v_{ni})$ is also a permutation of $(u_{1i},\cdots,u_{ni})$ with $i=1,\cdots,d.$ From Vieta’s theorem, note that $v_{1i},\cdots,v_{ni}$ determine the same $B_{1i},\cdots,B_{ni}$ in some neighborhood of $y,$ i.e., $B_{1i},\cdots,B_{ni}$ can continue analytically to $y.$ Hence, $(i)$ holds. If $n<\nu,$ then we are able to determine another set $\{B^{\prime}_{ji}\}_{i,j}$ using the rest mapping elements $\{(x,H_{j})\}_{j}.$ However, it implies that Eq. (6) is reducible, a contradiction. ∎
According to Remark 2.16 and Theorem 2.19, we obtain:
Corollary 2.20.
$\mathcal{M}^{0}_{F}$ is a path-connected space.
2.4. Properties of Mapping Elements near Multiple Points
Let $(x_{0},y_{0})\in M\times N$ be a pair such that
$\Psi_{i}(x_{0},c_{i})=0,\ \ \ \ \Psi_{i,i}(x_{0},c_{i})=0$
for some $i\in\{1,\cdots,d\},$ where
$c_{i}=\frac{\zeta_{i}\circ\jmath(y_{0})}{\zeta_{0}\circ\jmath(y_{0})}.$
$1^{\circ}$ $x_{0}\not\in\cup_{i=1}^{d}A^{-1}_{i\nu_{i}}(0)$
Choosing a sufficiently small neighborhood $U(x_{0})$ of $x_{0}$ such that $A_{i\nu_{i}}\not=0$ on $\overline{U(x_{0})}$ for $i=1,\cdots,d.$ According to the continuation of $A_{ij}$ for $1\leq i\leq d$ and $0\leq j\leq\nu_{i},$ there exists a number $M_{0}>0$ such that
(8) $0\leq\mathop{\mathrm{min}}_{\begin{subarray}{c}1\leq i\leq d\\ 0\leq j\leq\nu_{i}-1\end{subarray}}\left|\frac{A_{ij}}{A_{i\nu_{i}}}\right|\leq\max_{\begin{subarray}{c}1\leq i\leq d\\ 0\leq j\leq\nu_{i}-1\end{subarray}}\left|\frac{A_{ij}}{A_{i\nu_{i}}}\right|\leq M_{0},\ \ \ \ ^{\forall}x\in\overline{U(x_{0})}.$
According to Corollary 2.11, for any $x,y\in U(x_{0})\setminus\mathscr{M}_{F}$ with $x\not=y,$ there are two groups of distinct quasi-regular mapping elements $\{x,H_{j}\}_{j=1}^{\nu},\{y,G_{j}\}_{j=1}^{\nu}$ of $F.$ Applying Theorem 2.17 again, for a curve $\gamma:[0,1]\rightarrow U(x_{0})\setminus\mathscr{M}_{F}$ with $\gamma(0)=x$ and $\gamma(1)=y,$ one can extend $\{x,H_{j}\}_{j=1}^{\nu}$ analytically to $\{y,G_{j}\}_{j=1}^{\nu}$ along $\gamma.$ In other words, there exists a permutation $(i_{1},\cdots,i_{\nu})$ of $(1,\cdots,\nu)$ such that $H_{1},\cdots,H_{\nu}$ can be continued analytically to $G_{i_{1}},\cdots,G_{i_{\nu}}$ along $\gamma,$ respectively. The symmetric property implies that $\{y,G_{j}\}_{j=1}^{\nu}$ shall be also continued analytically to $\{x,H_{j}\}_{j=1}^{\nu}$ along $\gamma.$ However, we may mention that $G_{i_{1}},\cdots,G_{i_{\nu}}$ won’t be, in general, continued analytically back to $H_{1},\cdots,H_{\nu}$ along $\gamma,$ respectively. In what follows, we consider the analytic continuation of a mapping element along a closed Jordan curve surrounding any multiple point $x_{0}\in\mathscr{M}_{F}.$ Taking a smooth closed Jordan curve $\gamma_{0}\subset U(x_{0})$ around $x_{0}$ such that which is close to $x_{0}$ and doesn’t surround or cross any connected components of $\mathscr{M}_{F}$ that do not pass through $x_{0}.$ Without loss of generality, let us look at the analytic continuation of $H_{1}$ along $\gamma_{0}$ starting from $y_{0}\in\gamma_{0}.$ When $H_{1}$ returns to $y_{0}$ along $\gamma_{0},$ it will turn to a mapping $H_{k_{1}}\in\{H_{j}\}_{j=1}^{\nu},$ by which we mean that for a local coordinate chart $(U(x_{0}),\varphi)$ (here $U(x_{0})$ is small enough), $H_{1}\circ\varphi^{-1}(z)$ jumps to a single-valued component $H_{k_{1}}\circ\varphi^{-1}(z)$ when $z$ travels around $\varphi(x_{0})$ once along $\varphi(\gamma_{0})$ starting from $\varphi(y_{0}),$ due to the change of $m$-dimensional argument of $z=(z_{1},\cdots,z_{m}).$ If $k_{1}\not=1,$ then we will continue to extend $u_{k_{1}}$ analytically along $\gamma_{0}$ in the same direction of rotation. Repeat this action, we shall see finally that $H_{1}$ cycles on and on periodically. Let $\lambda_{1}$ be the smallest period such that $H_{1}$ can return to itself along all such curves $\gamma_{0}.$ Then, one can obtain $H_{1},H_{k_{1}},\cdots,H_{k_{\lambda_{1}-1}}.$ We call $x_{0}$ a branch point with order $\lambda_{1}-1$ of $H_{1}$ for $\lambda_{1}>1.$ Repeatedly, it can be seen that $\{H_{j}\}_{j=1}^{\nu}$ is divided into finitely many groups. Let $l$ be the number of groups. If the $j$-th group has $\lambda_{j}$ members for $j=1,\cdots,l,$ then
$\lambda_{1}+\cdots+\lambda_{l}=\nu.$
We just analyze the first group: $H_{1},H_{k_{1}},\cdots,H_{k_{\lambda_{1}-1}}.$ Take a suitable $U(x_{0}),$ there exists a biholomorphic mapping $\varphi:U(x_{0})\to\Delta^{m}(2)$ such that
$\varphi(x_{0})=\textbf{0},\ \ \ \ \varphi(\gamma_{0})=\big\{(0,\cdots,0,z_{m})\in\Delta^{m}(2):\ |z_{m}|=1\big\},$
where $\Delta^{m}(2)$ is the polydisc centered at 0 with polyradius $(2,\cdots,2)$ in $\mathbb{C}^{m}.$ Setting $S_{0}=\varphi^{-1}(\Delta_{m}(1)),$ where $\Delta_{m}(1)$ is the unit disc centered at 0 in the $z_{m}$-plane. For any $x\in S_{0}\cap\gamma_{0},$ we have
$H_{1}(z)=H_{1}\circ\varphi(x)=H_{1}(0,\cdots,0,z_{m}).$
By Corollary 2.10, $\zeta_{0}\circ\jmath\circ H_{1}$ has no zeros. Write $\jmath\circ H_{1}=[1:u_{11}:\cdots:u_{1d}],$ where
$u_{1i}=\frac{\zeta_{i}\circ\jmath\circ H_{1}}{\zeta_{0}\circ\jmath\circ H_{1}},\ \ \ \ i=1,\cdots,d$
are $\lambda_{1i}$-valued algebroid functions, respectively, with $\lambda_{1}=\lambda_{11}\cdots\lambda_{1d}.$ Set
$u^{*}_{1i}(\eta)=u_{1i}(0,\cdots,0,\eta^{\lambda_{1i}}),\ \ \ \ i=1,\cdots,d.$
We show that $u^{*}_{1i}$ is a single-valued analytic function on $\Delta_{m}(1).$ Note that $\lambda_{1}$ is the branch order of $x_{0},$ we deduce that $u^{*}_{1i}$ is holomorphic on $\Delta_{m}(1)\setminus\{\textbf{0}\},$ because $\varphi(\gamma_{0})$ is so close to 0 that there exist no any branch points of $u^{*}_{1i}$ in $\Delta_{m}(1)\setminus\{\textbf{0}\}.$ We only need to examine that $\zeta=0$ is a removed singularity of $u^{*}_{1i}.$ To do so, it suffices to show that $u^{*}_{1i}$ is bounded near 0 due to Riemann’s analytic continuation. In fact, it follows from (6) and (8) that
$\left|u_{1i}\right|$ $=$ $\frac{1}{\left|A_{i\nu_{i}}{u_{1i}}^{\nu_{i}-1}\right|}\left|A_{i(\nu_{i}-1)}{u_{1i}}^{\nu_{i}-1}+\cdots+A_{i0}\right|$
$\leq$ $M_{0}\left(1+\frac{1}{|u_{1i}|}+\cdots+\frac{1}{|u_{1i}|^{\nu_{i}-1}}\right)$
on $\overline{U(x_{0})}.$ If $|u_{1i}|\geq 1,$ then $|u_{1i}|\leq\nu_{i}M_{0}.$ Hence, it gives
$|u_{1i}|\leq\max\{1,\nu_{i}M_{0}\},\ \ \ \ ^{\forall}x\in\overline{U(x_{0})}.$
That is, $u_{1i}$ is bounded on $\overline{U(x_{0})}.$ It means that $u^{*}_{1i}$ is holomorphic on $\Delta_{m}(1).$ For $i=1,\cdots,d,$ we can expand $u^{*}_{1i}$ into the following Taylor series
(9) $u^{*}_{1i}(\eta)=b_{i0}+b_{i\tau_{i}}\eta^{\tau_{i}}+b_{i(\tau_{i}+1)}\eta^{\tau_{i}+1}+\cdots$
[…]
$\varphi(x_{0})=\textbf{0};\ \ \ \ \varphi(x)=(\hat{z}_{m}(x),0),\ \ \ \ ^{\forall}x\in\mathscr{B}_{H_{1}},$
such that each $u_{1i}$ can be expanded into the following Pusieux series
$u_{1i}(z)=B_{i0}(\hat{z}_{m})+B_{i\tau_{i}}(\hat{z}_{m})z_{m}^{\frac{\tau_{i}}{\lambda_{1}}}+B_{i(\tau_{i}+1)}(\hat{z}_{m})z_{m}^{\frac{\tau_{i}+1}{\lambda_{1}}}+\cdots$
at $x_{0},$ where $B_{i0},B_{i\tau_{i}},\cdots$ are holomorphic functions in $\hat{z}_{m}=(z_{1},\cdots,z_{m-1}).$ Comparing coefficients, we deduce that
$B_{ij}(\textbf{0})=b_{ij},\ \ \ \ j=0,\tau_{i},\tau_{i}+1,\cdots$
According to the above arguments, we show that $H_{1}$ is a $\lambda_{1}$-valued analytic mapping defined in a neighborhood of $x_{0}.$ We call $(x_{0},H_{1})$ a $\lambda_{1}$-leaf mapping element. If one cuts $U(x_{0})$ along a real hypersurface passing through $\mathscr{B}_{H_{1}},$ then $H_{1}$ can be separated into $\lambda_{1}$ single-valued analytic components, and all these $\lambda_{1}$ components can turn to each other when they continue analytically along a closed curve around $x_{0}$ in $U(x_{0}).$
$2^{\circ}$ $x_{0}\in\cup_{i=1}^{d}A^{-1}_{i\nu_{i}}(0)$
Assume that $x_{0}$ is not a singular branch point of $H_{1}.$ For $i=1,\cdots,d,$ we set $W_{i}=1/V_{i}$ and define
$\Phi_{i}(x,W_{i})$ $:=$ $W_{i}^{\nu_{i}}\Psi_{j}(x,V_{i})$
$=$ $A_{i\nu_{i}}(x)+A_{i(\nu_{i}-1)}(x)W_{i}+\cdots+A_{i0}(x)W^{\nu_{i}}_{i}.$
$a)$ $x_{0}\not\in\cup_{i=1}^{d}A^{-1}_{i0}(0)$
By the previous arguments, one can also expand $u_{1i}$ locally into a Pusieux series of the form
$u_{1i}(z)=C_{i0}(\hat{z}_{m})+C_{i\tau_{i}}(\hat{z}_{m})z_{m}^{\frac{\tau_{i}}{\lambda_{1}}}+C_{i(\tau_{i}+1)}(\hat{z}_{m})z_{m}^{\frac{\tau_{i}+1}{\lambda_{1}}}+\cdots$
at $x_{0}.$ If $C_{i0}\not\equiv 0,$ then $u_{1i}^{-1}$ can be expanded at $x_{0}$ in the form
$u_{1i}(z)^{-1}$ $=$ $\tilde{C}_{i(-\tau_{i})}(\hat{z}_{m})z_{m}^{-\frac{\tau_{i}}{\lambda_{1}}}+\tilde{C}_{i(-\tau_{i}+1)}(\hat{z}_{m})z_{m}^{-\frac{\tau_{i}-1}{\lambda_{1}}}+\cdots$
$+~\tilde{C}_{i(-1)}(\hat{z}_{m})z_{m}^{-\frac{1}{\lambda_{1}}}+\frac{1}{C_{i0}(\hat{z}_{m})}+\tilde{C}_{i1}(\hat{z}_{m})z_{m}^{\frac{1}{\lambda_{1}}}+\cdots$
$b)$ $x_{0}\in\cup_{i=1}^{d}A^{-1}_{i0}(0)$
Take a number $c_{0}$ with $\Psi_{i}(x_{0},c_{0})\not=0$ for all $i.$ Put $\tilde{W}_{i}=W_{i}-c_{0}.$ Then
$\tilde{\Psi}_{i}(x,\tilde{W}_{i})$ $:=$ $\Psi_{i}(x,\tilde{W}_{i}+c_{0})$
$=$ $\tilde{A}_{i\nu_{i}}(x)\tilde{W}_{i}^{\nu_{i}}+\tilde{A}_{i(\nu_{i}-1)}(x)\tilde{W}_{i}^{\nu_{i}-1}+\cdots+\tilde{A}_{i0}(x).$
Evidently, $\tilde{A}_{i0}(x_{0})=\tilde{\Psi}_{i}(x_{0},0)=\Psi_{i}(x_{0},c_{0})\not=0.$ Hence, the case $b)$ turns to the case $a).$
To conclude, for each $x\in M,$ there are $l=l(x)$ distinct multi-leaf mapping elements of $F,$ denoted by $(x,H_{1}),\cdots,(x,H_{l}),$ in which $H_{j}$ is a $\nu_{j}$-valued analytic mapping with $1\leq\nu_{j}\leq\nu$ and $\nu_{1}+\cdots+\nu_{l}=\nu.$ Note that $H_{j}$ has branch order $\nu_{j}-1$ at $x$ with $j=1,\cdots,l.$ If $l<\nu,$ then one says that $x$ is a branch point with order $\nu-l$ of $F.$ Let $\mathscr{B}_{F}$ denote the set of all branch points of $F,$ which is an analytic set in $M,$ called the branch set of $F.$ Recall that
$W_{i}=\frac{\zeta_{i}\circ\jmath\circ F}{\zeta_{0}\circ\jmath\circ F},\ \ \ \ i=1,\cdots,d.$
We have shown that
Theorem 2.21.
Let $(x_{0},H)$ be a $\lambda$-leaf mapping element of $F.$ Assume that
$u_{i}=\frac{\zeta_{i}\circ\jmath\circ H}{\zeta_{0}\circ\jmath\circ H},\ \ \ \ i=1\cdots,d$
are $\lambda_{i}$-valued components of $W_{i}$ in a neighborhood of $x_{0},$ respectively. Then, we have $\lambda=\lambda_{1}\cdots\lambda_{d}.$ Moreover, if $x_{0}$ is not a singular branch point of $H,$ then $u_{i}$ can be expanded into the following convergent Pusieux series
$u_{i}(z)=B_{i0}(\hat{z}_{m})+B_{i\tau_{i}}(\hat{z}_{m})z_{m}^{\frac{\tau_{i}}{\lambda_{i}}}+B_{i(\tau_{i}+1)}(\hat{z}_{m})z_{m}^{\frac{\tau_{i}+1}{\lambda_{i}}}+\cdots$
at $x_{0}$ in a local holomorphic coordinate $z=(z_{1},\cdots,z_{m})$ such that $z(x_{0})$=0 and $z(x)=(\hat{z}_{m}(x),0)$ for all $x\in\mathscr{B}_{H},$ where $\mathscr{B}_{H}$ denotes the branch set of $H$ in a neighborhood of $x_{0},$ and $B_{i0},B_{i\tau_{i}},\cdots$ are meromorphic functions in $\hat{z}_{m}=(z_{1},\cdots,z_{m-1}).$
Let $F_{1},\cdots,F_{\nu}$ be $\nu$ distinct single-valued components of $F.$ Assume that $(U(x_{0}),H)$ is a $\lambda$-leaf mapping element of $F$ at $x_{0}$ with distinct single-valued components $(U(x_{0}),H_{1}),\cdots,(U(x_{0}),H_{\lambda}).$ Then, there exists a permutation $(i_{1},\cdots,i_{\nu})$ of $(1,\cdots,\nu)$ such that $H_{1}=F_{1},\cdots,H_{\lambda}=F_{i_{\lambda}}$ on $U(x_{0}).$
3. Multi-leaf Complex Manifolds by Algebroid Mappings
Let $\tilde{x}=(U(x),H)\in\mathcal{M}^{0}_{F}$ be any mapping element. A neighborhood V(x~)V($\tilde${x}) of $\tilde{x}$ means a set of all $\tilde{y}\in\mathcal{M}^{0}_{F}$ with $y\in U(x),$ such that $\tilde{y}$ is a direct analytic continuation of $\tilde{x}$, i.e., $V(\tilde{x})=\{(y,H):y\in U(x)\}.$ For a proper $U(x),$ there is a biholomorphic mapping $\varphi_{x}:U(x)\to\mathbb{B}^{m}(1)$ such that $\varphi_{x}(x)=\textbf{0},$ where $\mathbb{B}^{m}(1)$ is the unit ball centered at 0 in $\mathbb{C}^{m}.$ Define a natural mapping
$\psi_{\tilde{x}}:V(\tilde{x})\to\mathbb{B}^{m}(1)$
via
[…]
Clearly, $\psi_{\tilde{x}}$ is a homeomorphism whose inverse is
$\psi^{-1}_{\tilde{x}}(z)=\big(\varphi^{-1}_{x}(z),H\big),\ \ \ \ ^{\forall}z\in\mathbb{B}^{m}(1).$
Set
$\mathscr{U}_{F}=\left\{(V(\tilde{x}),\psi_{\tilde{x}}):\tilde{x}\in\mathcal{M}^{0}_{F}\right\}.$
[…]
$\psi_{\tilde{y}}\circ\psi^{-1}_{\tilde{x}}(z)=\psi_{\tilde{y}}\big(\varphi^{-1}_{x}(z),H\big)=\psi_{\tilde{y}}\big(\varphi^{-1}_{x}(z),G\big)=\varphi_{y}\circ\varphi^{-1}_{x}(z)$
for any $).z\in\psi_{\tilde{x}}(V(\tilde{x})\cap V(\tilde{y}$)). Similarly, we have
$).\psi_{\tilde{x}}\circ\psi^{-1}_{\tilde{y}}(z)=\varphi_{x}\circ\varphi^{-1}_{y}(z),\ \ \ \ ^{\forall}z\in\psi_{\tilde{y}}(V(\tilde{x})\cap V(\tilde{y}$)).
It shows that $\psi_{\tilde{y}}\circ\psi^{-1}_{\tilde{x}}$ is a biholomorphic mapping. On the other hand, each $x\in M\setminus\mathscr{B}_{F}$ corresponds to $\nu$ distinct mapping elements $\{(x,H_{j})\}_{j=1}^{\nu}\subset\mathcal{M}^{0}_{F}.$ That is, we conclude that
Theorem 3.1.
$\mathcal{M}^{0}_{F}$ is a non-compact $\nu$-leaf complex manifold of complex dimension $m.$
Set
$\mathcal{M}^{1}_{F}=\big\{\text{all regular mapping elements of $F$}\big\}/\sim.$
Corollary 3.2.
$\mathcal{M}^{1}_{F}$ is a non-compact $\nu$-leaf complex manifold of complex dimension $m,$ which is a submanifold of $\mathcal{M}^{0}_{F}.$
More generally, we would construct a $\nu$-leaf complex manifold determined by all the mapping elements of $F.$ Let $\tilde{x}_{1}=(U(x_{1}),H_{1}),\tilde{x}_{2}=(U(x_{2}),H_{2})$ be any two mapping elements of $F.$ We define an equivalent relation as follows: $\tilde{x}_{1}\sim\tilde{x}_{2}$ if and only if
$x_{1}=x_{2};\ \ \ \ H_{10}(x)=H_{20}(x),\ \ \ \ ^{\forall}x\in\left(U(x_{1})\setminus\mathscr{B}_{H_{1}}\right)\cap\left(U(x_{2})\setminus\mathscr{B}_{H_{2}}\right)$
for some single-valued component $H_{j0}$ of $H_{j}$ with $j=1,2,$ where $\mathscr{B}_{H_{j}}$ is the branch set of $H_{j}$ for $j=1,2.$ Set
[…]
Definition 3.3.
A quasi-regular mapping element $(y,G)$ is called a direct analytic continuation of a multi-leaf mapping element $(U(x),H),$ if both $y\in U(x)\setminus\mathscr{B}_{H}$ and $G=H_{0}$ in a neighborhood of $y$ for some single-valued component $H_{0}$ of $H,$ where $\mathscr{B}_{H}$ is the branch set of $H.$ If so, one may regard that $(y,G)=(y,H_{0}).$
Let $\tilde{x}=(U(x),H)\in\mathcal{M}_{F}.$ A neighborhood V(x~)V($\tilde${x}) of $\tilde{x}$ is defined to be the set that contains the following elements:
  1. $\bullet$
    all $\tilde{y}\in\mathcal{M}_{F}$ with $y\in U_{1}(x)\setminus\mathscr{B}_{H}$ satisfying that $\tilde{y}$ is a direct analytic continuation of $\tilde{x}$;
  2. $\bullet$
    all $(y,H)\in\mathcal{M}_{F}$ with $y\in U_{1}(x)\cap\mathscr{B}_{H}.$ In the above, $U_{1}(x)\subset U(x)$ is a neighborhood of $x.$
In particular, a $r$-neighborhood $V(\tilde{x},\rho)$ of $\tilde{x}$ with $B(x,r)\subset U(x)$ is defined to be the set that contains the following elements:
  1. $\bullet$
    all $\tilde{y}\in\mathcal{M}_{F}$ with $y\in B(x,r)\setminus\mathscr{B}_{H}$ satisfying that $\tilde{y}$ is a direct analytic continuation of $\tilde{x}$;
  2. $\bullet$
    all $(y,H)\in\mathcal{M}_{F}$ with $y\in B(x,r)\cap\mathscr{B}_{H}.$
A subset $\mathcal{E}\subset\mathcal{M}_{F}$ is called an open set, if $\mathcal{E}$ is an empty set or each point $\tilde{x}\in\mathcal{E}$ is an inner point, i.e., there exists a $\epsilon$-neighborhood $V(\tilde{x},\epsilon)$ of $\tilde{x}$ such that $V(\tilde{x},\epsilon)\subset\mathcal{E}.$ It gives a topology of $\mathcal{M}_{F},$ which contains the topology of $\mathcal{M}^{0}_{F}.$ If $\tilde{x}$ is a $\lambda$-leaf mapping element, then V(x~)V($\tilde${x}) has $\lambda$ leaves at $\tilde{x}$. We call V(x~)V($\tilde${x}) a $\lambda$-leaf neighborhood of $\tilde{x}$.
Theorem 3.4.
$\mathcal{M}_{F}$ is a connected Hausdorff space.
Proof.
The Hausdorff property of $\mathcal{M}_{F}$ is proved using the similar arguments as in the proof of Theorem 5.11. We show that $\mathcal{M}_{F}$ is connected. Otherwise, we may assume that $\mathcal{M}_{F}=\mathcal{A}\cup\mathcal{B},$ where $\mathcal{A},\mathcal{B}$ are two disjoint open subsets of $\mathcal{M}_{F}.$ Set $\mathcal{S}=\{\tilde{x}\in\mathcal{M}_{F}:x\in\mathscr{B}_{F}\}.$ It is clear that $\mathcal{M}^{0}_{F}=(\mathcal{A}\setminus\mathcal{S})\cup(\mathcal{B}\setminus\mathcal{S}).$ Since $\mathscr{B}_{F}$ is an analytic set of complex codimension $1,$ we see that $\mathcal{A}\setminus\mathcal{S},\mathcal{B}\setminus\mathcal{S}$ are two disjoint open subsets of $\mathcal{M}^{0}_{F}.$ This implies that $\mathcal{M}^{0}_{F}$ is not connected, which is a contradiction. ∎
Next, we endow $\mathcal{M}_{F}$ with a complex structure such that $\mathcal{M}_{F}$ is a complex manifold. Let $\tilde{x}=(U(x),H)\in\mathcal{M}_{F}$ be any $\lambda$-leaf mapping element, in which $x$ is not a singular branch point of $H.$ Assume that
$u_{i}=\frac{\zeta_{i}\circ\jmath\circ H}{\zeta_{0}\circ\jmath\circ H},\ \ \ \ i=1,\cdots,d$
are $\lambda_{i}$-leaf function elements of $W_{i}$ on $U(x),$ respectively. By Theorem 2.21, we have $\lambda=\lambda_{1}\cdots\lambda_{d}.$ In further, choosing $U(x)$ properly if necessary, there exists a biholomorphic mapping $\varphi_{x}:U(x)\to\Delta^{m}(1)$ such that every $u_{i}$ can be expanded into the convergent Pusieux series
$u_{i}(z)$ $=$ $u_{i}\circ\varphi^{-1}_{x}(z)$
$=$ $B_{i0}(\hat{z}_{m})+B_{i\tau}(\hat{z}_{m})z_{m}^{\frac{\tau_{i}}{\lambda_{i}}}+B_{i(\tau_{i}+1)}(\hat{z}_{m})z_{m}^{\frac{\tau_{i}+1}{\lambda_{i}}}+\cdots$
on $U(x)$ with $\varphi_{x}(x)=\textbf{0}$ and $\varphi_{x}(y)=(\hat{z}_{m}(y),0)$ for all $y\in\mathscr{B}_{H}.$
Let V(x~)V($\tilde${x}) be a $\lambda$-leaf neighborhood of $\tilde{x}$ containing the following elements:
  1. $\bullet$
    all $\tilde{y}\in\mathcal{M}_{F}$ with $y\in U(x)\setminus\mathscr{B}_{H}$ satisfying that $\tilde{y}$ is a direct analytic continuation of $\tilde{x}$;
  2. $\bullet$
    all $(y,H)\in\mathcal{M}_{F}$ with $y\in U(x)\cap\mathscr{B}_{H}.$
Let $H_{1},\cdots,H_{\lambda}$ denote $\lambda$ distinct single-valued components of $H.$ Define a mapping $\psi_{\tilde{x}}:V(\tilde{x})\to\Delta^{m}(1)$ by
$\psi_{\tilde{x}}(y,H)$ $=$ $\big(\hat{z}_{m}(y),0\big),\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\ \ ^{\forall}y\in U(x)\cap\mathscr{B}_{H};$
$\psi_{\tilde{x}}(y,H_{j})$ $=$ $\Big(\hat{z}_{m}(y),\ |z_{m}(y)|^{\frac{1}{\lambda}}e^{i\frac{\arg z_{m}(y)+2(j-1)\pi}{\lambda}}\Big),\quad^{\forall}y\in U(x)\setminus\mathscr{B}_{H}$
for $j=1,\cdots,\lambda,$ where $z(y)=\varphi_{x}(y).$ We note that $\psi_{\tilde{x}}$ is a homeomorphism and its inverse is
$\psi^{-1}_{\tilde{x}}(z)=\begin{cases}\big(\varphi^{-1}_{x}(z),H\big),&{}^{\forall}z_{m}=0;\\ \left(\varphi_{x}^{-1}\big(\hat{z}_{m},|z_{m}|^{\lambda}e^{i\lambda\arg z_{m}}\big),H_{j}\right),&{}^{\forall}z_{m}\not=0,\end{cases}$
where $j$ is a positive integer such that $\big(2(j-1)\pi-\lambda\arg z_{m}\big)\in[0,2\pi).$ Denote by $\mathcal{M}_{F,\rm{sing}}$ the set of all mapping elements $(x,H)\in\mathcal{M}_{F}$ satisfying that $x$ is a singular branching point of $H,$ and by $\mathscr{B}_{F,\rm sing}$ the set of all corresponding singular branching points. Note that $\mathscr{B}_{F,\rm sing}$ is an analytic set of complex codimension not smaller than $2.$ Set
$\mathscr{U}=\big\{(V(\tilde{x}),\psi_{\tilde{x}}):\tilde{x}\in\mathcal{M}^{*}_{F}\big\},\ \ \ \ \mathcal{M}^{*}_{F}=\mathcal{M}_{F}\setminus\mathcal{M}_{F,\rm{sing}}.$
[…]
$)\psi_{\tilde{y}}\circ\psi^{-1}_{\tilde{x}}:\ \psi_{\tilde{x}}\big(V(\tilde{x})\cap V(\tilde{y})\big)\to\psi_{\tilde{y}}\big(V(\tilde{x})\cap V(\tilde{y}$$\big)$
[…]
$\psi_{\tilde{y}}\circ\psi^{-1}_{\tilde{x}}(z)$ $=$ $\psi_{\tilde{y}}\left(\varphi_{x}^{-1}\big(\hat{z}_{m},|z_{m}|^{\lambda}e^{i\lambda\arg z_{m}}\big),H_{j}\right)$
$=$ $\psi_{\tilde{y}}\left(\varphi_{x}^{-1}\big(\hat{z}_{m},|z_{m}|^{\lambda}e^{i\lambda\arg z_{m}}\big),G\right)$
$=$ $\varphi_{y}\circ\varphi^{-1}_{x}\Big(\hat{z}_{m},|z_{m}|^{\lambda}e^{i\lambda\arg z_{m}}\Big)$
$=$ $\varphi_{y}\circ\varphi^{-1}_{x}\big(\hat{z}_{m},z_{m}^{\lambda}\big),$
where $j$ is a positive integer such that $\big(2(j-1)\pi-\lambda\arg z_{m}\big)\in[0,2\pi).$ Hence, $\psi_{\tilde{y}}\circ\psi^{-1}_{\tilde{x}}$ is a holomorphic mapping. On the other hand, its inverse is
$\psi_{\tilde{x}}\circ\psi^{-1}_{\tilde{y}}(z)$ $=$ $\psi_{\tilde{x}}\big(\varphi_{y}^{-1}(z),G\big)$
$=$ $\psi_{\tilde{x}}\big(\varphi_{y}^{-1}(z),H_{j}\big)$
$=$ $\varphi_{x}\circ\varphi^{-1}_{y}\Big(\hat{z}_{m},\ |z_{m}|^{\frac{1}{\lambda}}e^{i\frac{\arg z_{m}+2(j-1)\pi}{\lambda}}\Big)$
$=$ $\varphi_{x}\circ\varphi^{-1}_{y}\Big(\hat{z}_{m},\big(z_{m}^{\frac{1}{\lambda}}\big)_{j}\Big),$
where $\big(z_{m}^{1/\lambda}\big)_{j}$ is the $j$-th single-valued component of $z_{m}^{1/\lambda}.$ Evidently, $\psi_{\tilde{x}}\circ\psi^{-1}_{\tilde{y}}$ is also a holomorphic mapping. Hence, $\mathcal{M}^{*}_{F}$ is a non-compact $\nu$-leaf complex manifold of complex dimension $m.$
We finally treat the singular branch set $\mathcal{M}_{F,\rm sing}.$ Let $\pi:\mathcal{M}_{F}\to M$ be the natural projection which sends any mapping element $(x,H)$ to $x.$ Note that $\pi:\mathcal{M}^{*}_{F}\to M\setminus\mathscr{B}_{F,\rm sing}$ is a $\nu$-sheeted ramified analytic covering such that
$M=\pi(\mathcal{M}^{*}_{F})\cup\pi(\mathcal{M}_{F,\rm sing})=\pi(\mathcal{M}^{*}_{F})\cup\mathscr{B}_{F,\rm{sing}}.$
Since $\mathscr{B}_{F,{\rm sing}}$ is an analytic set of complex codimension not smaller than $2,$ the complex structure of $M$ is a natural extension of the complex structure of $\pi(\mathcal{M}^{*}_{F})=M\setminus\mathscr{B}_{F,\rm sing}.$ It implies that we can naturally extend the complex structure of $\mathcal{M}^{*}_{F}$ to $\mathcal{M}_{F,\rm sing}$ via $\pi.$
Therefore, we have concluded that
Theorem 3.5.
$\mathcal{M}_{F}$ is a $\nu$-sheeted ramified analytic covering of $M.$ We call $\mathcal{M}_{F}$ the $\nu$-leaf complex manifold by $F.$
It is clear that the above theorem implies that
Corollary 3.6 (Uniformization).
An algebroid mapping $F:M\to N$ can be lifted to a meromorphic mapping $f:\mathcal{M}_{F}\to N$ via the natural projection $\pi:\mathcal{M}_{F}\to M$ by mapping $(x,H)$ to $x$ such that $F=f\circ\pi^{-1}.$
4. Estimate of Branch Divisors
[…]
We need the following lemma:
Lemma 4.1.
For $i=1,\cdots,d,$ we have
$\left(W_{i}=\infty\right)=(W_{i1}=\infty)+\cdots+(W_{i\nu_{i}}=\infty)=(A_{i\nu_{i}}=0).$
Proof.
Let $x_{0}$ be an arbitrary zero of $A_{i\nu_{i}}.$ Since the indeterminacy locus $I_{\mathscr{A}_{i}}$ of
$\mathscr{A}_{i}=[A_{i0}:\cdots:A_{i\nu_{i}}]:\ M\rightarrow\mathbb{P}^{\nu_{i}}(\mathbb{C})$
is either empty or of complex codimension 2, we obtain $I_{\mathscr{A}_{i}}=0$ in the sense of divisors. Thus, we may assume that $x_{0}\not\in I_{\mathscr{A}_{i}}.$ It follows that
(10) $n\big(x_{0},A_{i\nu_{i}}=0\big)=\max_{0\leq j\leq\nu_{i}-1}n\Big(x_{0},\frac{A_{ij}}{A_{i\nu_{i}}}=\infty\Big),$
where $n(x_{0},g=a)$ denotes the number of $a$-valued points of a function $g$ at $x_{0}.$ Without loss of generality, we assume that $W_{i1},\cdots,W_{ip}$ are all those components that take $\infty$ at $x_{0},$ According to Vieta’s theorem, we obtain
$\sum_{1\leq j\leq\nu_{i}}W_{ij}$ $=$ $-\frac{A_{i(\nu_{i}-1)}}{A_{\nu_{i}}}$
$\cdots\cdots\cdots$
$\sum_{1\leq j_{1}<\cdots<j_{p}\leq\nu_{i}}W_{ij_{1}}\cdots W_{ij_{p}}$ $=$ $(-1)^{p}\frac{A_{i(\nu_{i}-p)}}{A_{i\nu_{i}}}$
$\cdots\cdots\cdots$
$W_{i1}\cdots W_{i\nu_{i}}$ $=$ $(-1)^{\nu_{i}}\frac{A_{i0}}{A_{i\nu_{i}}}$
which implies that $A_{i(\nu_{i}-p)}(x_{0})\not=0,$ because $w_{i1}\cdots w_{ip}$ reaches the maximal pole order at $x_{0}\not\in I_{\mathscr{A}_{i}}.$ It yields from (10) that
$n\big(x_{0},W_{i}=\infty\big)$
$=$ $n(x_{0},W_{i1}=\infty)+\cdots+n(x_{0},W_{ip}=\infty)$
$=$ $n(x_{0},W_{i1}\cdots W_{ip}=\infty)$
$=$ $n\Big(x_{0},\frac{A_{i(\nu_{i}-p)}}{A_{\nu_{i}}}=\infty\Big)$
$=$ $n\big(x_{0},A_{i\nu_{i}}=0\big),$
which leads to that $(W_{i}=\infty)=(A_{i\nu_{i}}=0).$
Setting $V_{i}=1/W_{i}$ and using the similar arguments, we can also show the following corollary.
Corollary 4.2.
For $i=1,\cdots,d,$ we have
$\left(W_{i}=0\right)=(W_{i1}=0)+\cdots+(W_{i\nu_{i}}=0)=(A_{i0}=0).$
For any $x\in\mathscr{B}_{F},$ we denote by $(x,H_{1}),\cdots,(x,H_{l})$ the all $l=l(x)$ distinct mapping elements of $F$ at $x$ with $\lambda_{1},\cdots,\lambda_{l}$ leaves, respectively. Recall that $\lambda_{1}+\cdots+\lambda_{l}=\nu$ and $x$ is a $(\nu-l)$-order branching point of $F.$ The branching divisor of $F$ is a divisor $\mathscr{D}_{F}$ on $M$ such that
${\rm{Supp}}\mathscr{D}_{F}=\mathscr{B}_{F},\ \ \ \ {\rm{Ord}}_{x}\mathscr{D}_{F}=\nu-l(x),\ \ \ \ ^{\forall}x\in\mathscr{B}_{F}.$
Let $(x_{0},H)$ be any $\lambda$-leaf mapping element of $F$ at $x_{0}.$ One may assume without loss of generality that $\zeta_{0}\circ\jmath\circ H\not\equiv 0.$ Assume that
$u_{i}=\frac{\zeta_{i}\circ\jmath\circ H}{\zeta_{0}\circ\jmath\circ H},\ \ \ \ i=1,\cdots,d$
are $\lambda_{i}$-valued components of $W_{i}$ at $x_{0},$ respectively, i.e., $(x_{0},u_{i})$ are $\mu_{i}$-leaf function elements of $W_{i}$ at $x_{0},$ respectively. Note that
(11) $\mu_{1}\cdots\mu_{d}=\lambda.$
Write $u_{i}=\{u_{ij}\}_{j=1}^{\mu_{i}}$ for $i=1,\cdots,d.$ Then, there are $\mu_{i}$ distinct single-valued components of $W_{i},$ saying without loss of generality that $W_{i1},\cdots,W_{i\mu_{i}},$ such that $u_{ij}=W_{ij}$ for $j=1,\cdots,\mu_{i}$ in some neighborhood of $x_{0}.$ Let $F_{1},\cdots,F_{\nu}$ be the $\nu$ distinct single-valued components of $F.$ Then
$\jmath\circ F_{i}:=\big[f_{i0}:\cdots:f_{id}\big],\ \ \ \ i=1,\cdots,\nu$
are $\nu$ distinct single-valued components of $\jmath\circ F.$ For $i=1,\cdots,\nu,$ put
$g_{ij}=\frac{f_{ij}}{f_{i0}},\ \ \ \ j=1,\cdots,d.$
It is clear that each $g_{ik}$ is a single-valued component of $W_{k}$ with $k=1,\cdots,d.$ By (11), it is not hard to deduce that
(12) $\Big\{\underbrace{W_{i1},\cdots,W_{i1}}_{\lambda/\mu_{i}};\cdots;\underbrace{W_{id},\cdots,W_{id}}_{\lambda/\mu_{i}}\Big\}\subset\big\{g_{1k},\cdots,g_{\nu k}\big\}.$
That is, there are $\lambda/\mu_{i}$ copies of $W_{ij}$ among $g_{1k},\cdots,g_{\nu k}$ for $j=1,\cdots,\mu_{i}.$
Set
(13) $J_{k}=\big(f_{10}\cdots f_{\nu 0}\big)^{2\nu-2}\prod_{1\leq i<j\leq\nu}\big(g_{ik}-g_{jk}\big)^{2},\ \ \ \ k=1,\cdots,d.$
Theorem 4.3.
For $k=1,\cdots,d,$ we have
$\mathscr{D}_{F}\leq\big(J_{k}=0\big).$
Proof.
For any $x_{0}\in\mathscr{B}_{F},$ we denote by $(x_{0},H_{1}),\cdots,(x_{0},H_{l})$ the all distinct mapping elements of $F$ at $x_{0},$ with $\lambda_{1},\cdots,\lambda_{l}$ leaves, respectively. Note that $\lambda_{1}+\cdots+\lambda_{l}=\nu$ as well as
(14) ${\rm{Ord}}_{x_{0}}\mathscr{D}=\sum_{s=1}^{l}(\lambda_{s}-1)=\nu-l.$
Assume that for $s=1,\cdots,l$
$u^{s}_{i}=\frac{\zeta_{i}\circ\jmath\circ H_{s}}{\zeta_{0}\circ\jmath\circ H_{s}},\ \ \ \ i=1,\cdots,d$
are $\mu^{s}_{i}$-valued components of $W_{i}$ at $x_{0},$ respectively, i.e., $(x_{0},u^{s}_{i})$ are $\mu^{s}_{i}$-leaf function elements of $W_{i}$ at $x_{0},$ respectively. Note that
$\mu^{s}_{1}\cdots\mu^{s}_{d}=\lambda_{s},\ \ \ \ s=1,\cdots,l.$
We write $u^{s}_{i}=\{u^{s}_{ij}\}_{j=1}^{\mu^{s}_{i}}$ for all $s,i.$ Then, there exist $\mu^{s}_{i}$ distinct single-valued components of $W_{i},$ saying without loss of generality that $W_{i1},\cdots,W_{i\mu^{s}_{i}},$ such that $u^{s}_{ij}=W_{ij}$ for $j=1,\cdots,\mu^{s}_{i}$ in some neighborhood of $x_{0}.$ Since the set of all singular branch points of $F$ has complex codimension not smaller than $2,$ it is sufficient to deal with the case where $x_{0}$ is a non-singular branch point of $F.$ By Theorem 2.21, we may write for all $s,i$
(15) $u^{s}_{i}(z)=B^{s}_{i0}(\hat{z}_{m})+B^{s}_{i\tau^{s}_{i}}(\hat{z}_{m})z_{m}^{\frac{\tau^{s}_{i}}{\mu^{s}_{i}}}+B^{s}_{i(\tau^{s}_{i}+1)}(\hat{z}_{m})z_{m}^{\frac{\tau^{s}_{i}+1}{\mu^{s}_{i}}}+\cdots$
in a local holomorphic coordinate $z=(\hat{z}_{m},z_{m})$ around $x_{0},$ such that $x_{0}$ has a local holomorphic coordinate $z=\textbf{0}.$
Take any $s\in\{1,\cdots,l\}.$ If $x_{0}$ is not a pole of $u^{s}_{i},$ then we have $B^{s}_{i0}(\textbf{0})\not$$\infty$. Using (12) and (15), we deduce that there exist at least $\lambda_{s}(\lambda_{s}-1)/2$ terms which contain the factor $z_{m}^{2/\mu^{s}_{i}}$ in $J_{k}.$ Since $\mu^{s}_{i}|\lambda_{s},$ there exist at least $\lambda_{s}-1$ terms which contain the factor $z_{m}$ in $J_{k}.$ If $x_{0}$ is a pole of $u^{s}_{i},$ then it is also a pole of $W_{i1}\cdots W_{i\mu^{s}_{i}}.$ According to Lemma 4.1 and (12), we are led to
${\rm{Ord}}_{x_{0}}\big(u^{s}_{i}=\infty\big)\leq{\rm{Ord}}_{x_{0}}\big(g_{1k}\cdots g_{\nu k}=\infty\big)={\rm{Ord}}_{x_{0}}\big(f_{10}\cdots f_{\nu 0}=0\big).$
Combining the both cases and applying (14), we conclude that
${\rm{Ord}}_{x_{0}}\mathscr{D}_{F}\leq{\rm{Ord}}_{x_{0}}\big(J_{k}=0\big).$
This completes the proof. ∎
5. Equidistribution of Meromorphic Mappings
Let $(M,g)$ be an $m$-dimensional complete non-compact Kähler manifold with Laplace-Beltrami operator $\Delta$ and Kähler form
$\alpha=\frac{\sqrt{-1}}{\pi}\sum_{i,j=1}^{m}g_{i\bar{j}}dz_{i}\wedge d\bar{z}_{j}.$
Given a family $\{\Delta(r)\}_{r>0}$ of precompact domains $\Delta(r)\subset M$ with smooth boundaries $\partial\Delta(r)$ which can exhaust $M.$ Fix a reference point $o\in\Delta(r).$ We use $g_{r}(o,x)$ to stand for the Green function of $\Delta/2$ for $\Delta(r)$ with a pole at $o$ satisfying Dirichlet boundary condition, i.e.,
$\frac{1}{2}\Delta g_{r}(o,x)=\delta_{o}(x),\ \ \ \ ^{\forall}x\in\Delta(r);\ \ \ \ \ g_{r}(o,x)=0,\ \ \ \ ^{\forall}x\in\partial\Delta(r),$
where $\delta_{o}$ is the Dirac’s delta function on $M$ with a pole at $o.$ Note that the harmonic measure $\pi_{r}$ on $\partial\Delta(r)$ with respect to $o$ is defined by
$d\pi_{r}(x)=\frac{1}{2}\frac{\partial g_{r}(o,x)}{\partial\vec{\nu}}d\sigma_{r}(x),$
where $d\sigma_{r}$ is the induced Riemannian volume element of $\partial\Delta(r)$ and $\partial/\partial\vec{\nu}$ is the inward normal derivative on $\partial\Delta(r).$ Moreover, let $(N,\omega)$ be a compact Kähler manifold of complex dimension $n.$
5.1. First Main Theorem
We introduce Nevanlinna’s functions. Let $f:M\to N$ be a meromorphic mapping. The characteristic function of $f$ with respect to $\omega$ is defined by
$T_{f}(r,\omega)$ $=$ $\frac{\pi^{m}}{(m-1)!}\int_{\Delta(r)}g_{r}(o,x)f^{*}\omega\wedge\alpha^{m-1}$
Since $N$ is compact, for another Kähler metric $\omega^{\prime}$ on $N,$ we have
$T_{f}(r,\omega^{\prime})=T_{f}(r,\omega)+O(1).$
Let $D$ be an effective divisor cohomologous to $\omega,$ by which we mean that there exists a function $u_{D}\geq 0$ on $N$ such that
(16) $\omega-[D]=2dd^{c}u_{D}$
in the sense of currents, here $d=\partial+\overline{\partial}$ and $d^{c}=(\sqrt{-1}/4\pi)(\overline{\partial}-\partial).$ Thus, $u_{D}$ is quasi-plurisubharmonic, i.e., $dd^{c}(-u_{D})\geq-\omega/2.$ The proximity function of $f$ with respect to $D$ is defined by
$m_{f}(r,D)=\int_{\partial\Delta(r)}u_{D}\circ fd\pi_{r}.$
Define the counting function of $f$ with respect to $D$ by
$N_{f}(r,D)$ $=$ $\frac{\pi^{m}}{(m-1)!}\int_{f^{*}D\cap\Delta(r)}g_{r}(o,x)\alpha^{m-1}.$
Moreover, the simple counting function of $f$ with respect to $D$ is defined by
$\overline{N}_{f}(r,D)$ $=$ $\frac{\pi^{m}}{(m-1)!}\int_{f^{-1}(D)\cap\Delta(r)}g_{r}(o,x)\alpha^{m-1}.$
Lemma 5.1 (Green-Dynkin Formula, [1, 5, 6]).
Let $u$ be a $\mathscr{C}^{2}$-class function on $M$ outside a polar set of singularities at most. Assume that $u(o)\not$$\infty$. Then
$\int_{\partial\Delta(r)}u(x)d\pi_{r}(x)-u(o)=\frac{1}{2}\int_{\Delta(r)}g_{r}(o,x)\Delta u(x)dv(x).$
We have the first main theorem as follows.
Theorem 5.2 (First Main Theorem).
Assume that $f(o)\not\in{\rm Supp}D.$ Then
$T_{f}(r,\omega)+u_{D}\circ f(o)=m_{f}(r,D)+N_{f}(r,D).$
Proof.
Let $\mathscr{O}(D)$ be the holomorphic line bundle over $N$ defined by $D.$ Let $s_{D}$ be the canonical section associated to $D,$ i.e., $s_{D}$ is a holomorphic section of $\mathscr{O}(D)$ over $N$ with zero divisor $D.$ Locally, we write $s_{D}=\tilde{s}_{D}e,$ where $e$ is a local holomorphic frame of $\mathscr{O}(D).$ By Poincaré-Lelong formula, we obtain $D]=dd^{c}[\log|\tilde{s}_{D}|^{2}]>0$ in the sense of currents. Furthermore, it yields from (16) that
$\omega=dd^{c}\big[2u_{D}+\log|\tilde{s}_{D}|^{2}\big]>0,$
which leads to that $2u_{D}+\log|\tilde{s}_{D}|^{2}$ is plurisubharmonic. Since $N$ is a Kähler manifold, we deduce that both $\log|\tilde{s}_{D}|^{2}$ and $2u_{D}+\log|\tilde{s}_{D}|^{2}$ are subharmonic. Thus, $u_{D}$ is the difference of two subharmonic functions. It implies that the set of poles of $u_{D}\circ f$ is polar. So, applying Green-Dynkin formula to $u_{D}\circ f,$ we obtain
$\int_{\partial\Delta(r)}u_{D}\circ fd\pi_{r}-u_{D}\circ f(o)=\frac{1}{2}\int_{\Delta(r)}g_{r}(o,x)\Delta(u_{D}\circ f)dv.$
The first term on the left hand side equals $m_{f}(r,D)$ and the right hand side is that
$\frac{1}{2}\int_{\Delta(r)}g_{r}(o,x)\Delta(u_{D}\circ f)dv$
$=$ $\frac{2\pi^{m}}{(m-1)!}\int_{\Delta(r)}g_{r}(o,x)dd^{c}(u_{D}\circ f)\wedge\alpha^{m-1}$
$=$ $\frac{\pi^{m}}{(m-1)!}\int_{\Delta(r)}g_{r}(o,x)f^{*}\omega\wedge\alpha^{m-1}$
$-\frac{\pi^{m}}{(m-1)!}\int_{\Delta(r)}g_{r}(o,x)dd^{c}\left[\log|\tilde{s}_{D}\circ f|^{2}\right]\wedge\alpha^{m-1}$
$=$ $T_{f}(r,\omega)-N_{f}(r,D).$
This completes the proof. ∎
5.2. Second Main Theorem
We work on two typical types of Kähler manifolds with either non-positive sectional curvature or non-negative Ricci curvature. On such manifolds, we would like to establish the second main theorem of meromorphic mappings.
A. Non-positively Curved Case
Let $(M,g)$ be a complete non-compact Kähler manifold with non-positive sectional curvature. Recall that $\kappa(r)$ stands for the lower bound of the Ricci curvature of $M$ on $B(r),$ defined by (3) in Introduction. In this part, we set $\Delta(r)=B(r)$ with $r>0.$
Put
(17) $\chi(s,t)=\left\{\begin{array}[]{ll}t,&s=0;\\ \frac{\sinh st}{s},&s\not=0.\end{array}\right$.
Consider the following Jacobi equation on $0,\infty)$:
(18) $G^{\prime\prime}+\kappa(t)G=0;\ \ \ \ G(0)=0,\ \ \ \ G^{\prime}(0)=1.$
Note that this equation has a unique continuous solution.
We give two-sided estimates of $G(t)$ as follows.
Lemma 5.3.
For $t\geq 0,$ we have
$\chi(0,t)\leq G(t)\leq\chi\left(\sqrt{-\kappa(t)},t\right),$
[…]
Proof.
Treat the following Jacobi equation on $0,\infty)$:
$H_{1}^{\prime\prime}=0;\ \ \ \ H_{1}(0)=0,\ \ \ \ H_{1}^{\prime}(0)=1,$
which is uniquely solved by $H_{1}(t)=t.$ Since $\kappa(t)\leq 0$ for $t\geq 0,$ the standard comparison argument in ODE theory shows that
$G(t)\geq H_{1}(t)=t=\chi(0,t).$
Fix any number $t_{0}>0.$ It is clear that $\kappa(t)\geq\kappa(t_{0})$ for $0\leq t\leq t_{0}.$ Treat the following initial value problem on $0,t_{0}]$:
$H_{2}^{\prime\prime}+\kappa(t_{0})H_{2}=0;\ \ \ \ H_{2}(0)=0,\ \ \ \ H_{2}^{\prime}(0)=1,$
which gives a unique solution $H_{2}(t)=\chi(\sqrt{-\kappa(t_{0})},t).$ Applying the standard comparison argument (see [15], Theorems 1, 2 for instance), we deduce that
$G(t)\leq H_{2}(t)=\chi\left(\sqrt{-\kappa(t_{0})},t\right)$
for $0\leq t\leq t_{0}.$ In particular, $G(t_{0})\leq\chi(\sqrt{-\kappa(t_{0})},t_{0}).$ Using the arbitrariness of $t_{0},$ we deduce that
$G(t)\leq\chi\left(\sqrt{-\kappa(t)},t\right)$
for $t\geq 0.$ This completes the proof. ∎
Lemma 5.4 ([1]).
Let $\eta>0$ be a constant. Then, there exists a constant $c_{1}>0$ such that
$g_{r}(o,x)\geq c_{1}\frac{\displaystyle\int_{\rho(x)}^{r}G(t)^{1-2m}dt}{\displaystyle\int_{\eta}^{r}G(t)^{1-2m}dt}$
holds for all $x\in B(r)\setminus\overline{B(2\eta)}$ with $r>3\eta.$ In particular, if $M$ is non-parabolic, then there exists a constant $c_{2}>0$ such that
$g_{r}(o,x)\geq c_{2}\int_{\rho(x)}^{r}G(t)^{1-2m}dt$
[…]
For the upper estimate of $g_{r}(o,x),$ we have the following well-known fact.
Lemma 5.5 ([4]).
We have
[…]
and
$d\pi_{r}(x)\leq\frac{r^{1-2m}}{\omega_{2m-1}}d\sigma_{r}(x),\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \$
where $\omega_{2m-1}$ is the standard Euclidean area of the unit sphere in $\mathbb{R}^{2m}.$
We need the following Borel’s growth lemma (see [12, 17]):
Lemma 5.6.
Let $u\geq 0$ be a non-decreasing function on $(r_{0},\infty)$ with $r_{0}\geq 0.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(r_{0},\infty)$ of finite Lebesgue measure such that
$u^{\prime}(r)\leq u(r)^{1+\delta}$
holds for all $r>r_{0}$ outside $E_{\delta}.$
Proof.
The conclusion is clearly true for $u\equiv 0.$ Next, we assume that $u\not\equiv 0.$ Since $u\geq 0$ is a non-decreasing function, there exists a number $r_{1}>r_{0}$ such that $u(r_{1})>0.$ The non-decreasing property of $u$ implies that the limit $\eta:=\lim_{r\to\infty}u(r)$ exists or $\eta=\infty.$ If $\eta=\infty,$ then $\eta^{-1}=0.$ Set
$E_{\delta}=\left\{r\in(r_{0},\infty):\ u^{\prime}(r)>u(r)^{1+\delta}\right\}.$
Since $u$ is a non-decreasing function on $(r_{0},\infty),$ we deduce that $u^{\prime}(r)$ exists for almost all $r\in(r_{0},\infty).$ It is therefore
$\int_{E_{\delta}}dr$ $\leq$ $\int_{r_{0}}^{r_{1}}dr+\int_{r_{1}}^{\infty}\frac{u^{\prime}(r)}{u(r)^{1+\delta}}dr$
$=$ $\frac{1}{\delta u(r_{1})^{\delta}}-\frac{1}{\delta\eta^{\delta}}+r_{1}-r_{0}$
$<$ $\infty.$
This completes the proof. ∎
Set
(19) $K(r,\delta)=\frac{r^{1-2m}\bigg(\displaystyle\int_{\frac{1}{3}}^{r}G(t)^{1-2m}dt\bigg)^{(1+\delta)^{2}}}{G(r)^{(1-2m)(1+\delta)}},$
where $G(t)$ is defined by (18).
With the previous preparations, we establish a calculus lemma:
Theorem 5.7.
Let $k\geq 0$ be a locally integrable function on $M.$ Assume that $k$ is locally bounded at $o.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$\int_{\partial B(r)}kd\pi_{r}\leq CK(\delta,r)\bigg(\int_{B(r)}g_{r}(o,x)kdv\bigg)^{(1+\delta)^{2}}$
[…]
Proof.
For $r>1,$ it yields from Lemma 5.5 that
$\Lambda(r)$ $:=$ $\int_{B(r)}g_{r}(o,x)kdv$
$\geq$ $\int_{1}^{r}dt\int_{\partial B(t)}g_{r}(o,x)kd\sigma_{t}$
$\geq$ $\omega_{2m-1}\int_{1}^{r}t^{2m-1}dt\int_{\partial B(t)}g_{r}(o,x)kd\pi_{t}.$
According to Lemma 5.4, there exists a constant $c>0$ such that
$g_{r}(o,x)\geq c\frac{\displaystyle\int_{\rho(x)}^{r}G(t)^{1-2m}dt}{\displaystyle\int_{1/3}^{r}G(t)^{1-2m}dt}$
holds all $x\in B(r)\setminus\overline{B(2/3)}$ with $r>1.$ Whence, for $r>t>1$
$g_{r}(o,x)\big|_{\partial B(t)}\geq c\frac{\displaystyle\int_{t}^{r}G(s)^{1-2m}ds}{\displaystyle\int_{\frac{1}{3}}^{r}G(s)^{1-2m}ds}.$
It is therefore
$\int_{\partial B(t)}g_{r}(o,x)kd\pi_{t}$
$\geq$ $c\bigg(\displaystyle\int_{\frac{1}{3}}^{r}G(s)^{1-2m}ds\bigg)^{-1}\int_{\partial B(t)}kd\pi_{t}\displaystyle\int_{t}^{r}G(s)^{1-2m}ds.$
Combining the above to get
$\Gamma(r)$
$\geq$ $c\omega_{2m-1}\bigg(\displaystyle\int_{\frac{1}{3}}^{r}G(s)^{1-2m}ds\bigg)^{-1}\int_{1}^{r}t^{2m-1}dt\int_{\partial B(t)}kd\pi_{t}\displaystyle\int_{t}^{r}G(s)^{1-2m}ds$
$:=$ $\Lambda(r),$
which leads to
$\frac{\displaystyle\frac{d}{dr}\bigg(\Lambda(r)\displaystyle\int_{\frac{1}{3}}^{r}G(t)^{1-2m}dt\bigg)}{G(r)^{1-2m}}$ $=$ $c\omega_{2m-1}\int_{1}^{r}t^{2m-1}dt\int_{\partial B(t)}kd\pi_{t}.$
In further, we obtain
$\frac{d}{dr}\frac{\displaystyle\frac{d}{dr}\bigg(\Lambda(r)\displaystyle\int_{\frac{1}{3}}^{r}G(t)^{1-2m}dt\bigg)}{G(r)^{1-2m}}=c\omega_{2m-1}r^{2m-1}\int_{\partial B(r)}kd\pi_{r}.$
Using Borel’s growth lemma twice, then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$\int_{\partial B(r)}kd\pi_{r}$ $\leq$ $C\frac{\displaystyle r^{1-2m}\bigg(\displaystyle\int_{\frac{1}{3}}^{r}G(t)^{1-2m}dt\bigg)^{(1+\delta)^{2}}}{G(r)^{(1-2m)(1+\delta)}}\Lambda(r)^{(1+\delta)^{2}}$
$\leq$ $CK(r,\delta)\Gamma(r)^{(1+\delta)^{2}}$
holds for all $r>1$ outside $E_{\delta},$ where $C=1/c\omega_{2m-1}>0$ is clearly a constant independent of $k,\delta,r.$ This completes the proof. ∎
By estimating $\log^{+}K(r,\delta),$ we further obtain:
Corollary 5.8.
Let $k\geq 0$ be a locally integrable function on $M.$ Assume that $k$ is locally bounded at $o.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$\log^{+}\int_{\partial B(r)}kd\pi_{r}$
$\leq$ $(1+\delta)^{2}\log^{+}\int_{B(r)}g_{r}(o,x)kdv+O\left(\sqrt{-\kappa(r)}r+\delta\log r\right)$
[…]
Proof.
By Lemma 5.3, we have $G(t)\geq t$ for $t>0.$ Thus, one has
$\int_{\frac{1}{3}}^{r}G(t)^{1-2m}dt$ $\leq$ $\int_{1}^{r}t^{1-2m}dt+O(1)$
$\leq$ $\log r+O(1)$
for $r>1.$ It yields that
(20) $\log^{+}\left(\int_{\frac{1}{3}}^{r}G(t)^{1-2m}dt\right)^{(1+\delta)^{2}}\leq(1+\delta)^{2}\log^{+}\log r+O(1).$
If $\kappa\equiv 0,$ then $G(r)\leq r$ due to Lemma 5.3. Thus, we obtain
$r^{1-2m}G(r)^{(2m-1)(1+\delta)}\leq r^{(2m-1)\delta},$
which leads to
(21) $\log^{+}\left(r^{1-2m}G(r)^{(2m-1)(1+\delta)}\right)\leq(2m-1)\delta\log r$
for $r>1.$ Combining (20) with (21), we conclude that
$\log^{+}K(r,\delta)\leq O(\delta\log r).$
By Theorem 5.7, we have the conclusion holds. If $\kappa\not\equiv 0,$ then Lemma 5.3 gives that
$\log^{+}\left(r^{1-2m}G(r)^{(2m-1)(1+\delta)}\right)$
$\leq$ $\log^{+}G(r)^{(2m-1)(1+\delta)}$
$\leq$ $O\left(\sqrt{-\kappa(r)}r+1\right).$
Combining (20) with (5.2), we conclude that
$\log^{+}K(r,\delta)\leq O\left(\sqrt{-\kappa(r)}r+1\right).$
Hence, the conclusion also holds due to Theorem 5.7. ∎
Let $D_{1},\cdots,D_{q}$ be effective divisors on $N$ such that each $D_{j}$ is cohomologous to $\omega,$ i.e., there are functions $u_{D_{1}},\cdots,u_{D_{q}}\geq 0$ on $N$ such that
$\omega-[D_{j}]=2dd^{c}u_{D_{j}},\ \ \ \ j=1,\cdots,q.$
Set
$\Psi=\frac{\omega^{n}}{\prod_{j=1}^{q}u_{D_{j}}^{-2}e^{-2u_{D_{j}}}}.$
Lemma 5.9.
[2] Assume that $D_{1},\cdots,D_{q}$ are in general position. If $qw-{\rm Ric}\Psi>0,$ then
$(a)$ ${\rm Ric}\Psi\geq 0;$
$(b)$ there exists a constant $c>0$ such that $(-{\rm Ric}\Psi)^{n}\geq c\Psi;$
$(c)$ $\int_{N\setminus D}(-{\rm Ric}\Psi)^{n}<\infty,$ where $D=D_{1}+\cdots+D_{q}.$
Theorem 5.10.
Let $M$ be a complete non-compact Kähler manifold with non-positive sectional curvature. Let $(N,\omega)$ be a compact Kähler manifold of complex dimension not greater than that of $M.$ Let $D_{1},\cdots,D_{q}$ be effective divisors in general position on $N$ such that each $D_{j}$ is cohomologous to $\omega.$ Let $f:M\rightarrow N$ be a differentiably non-degenerate meromorphic mapping. Assume that $q\omega-{\rm Ric}(\omega^{n})>0.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$qT_{f}(r,\omega)+T_{f}(r,K_{N})+T(r,\mathscr{R})$
$\leq$ $\sum_{j=1}^{q}\overline{N}_{f}(r,D_{j})+O\big(\log T_{f}(r,\omega)+\sqrt{-\kappa(r)}r+\delta\log r\big)$
[…]
Proof.
Define a non-negative function $\xi$ by
$f^{*}\Psi\wedge\alpha^{m-n}=\xi\alpha^{m}.$
Using Poincaré-Lelong formula, we obtain
$dd^{c}\left[\log\xi\right]$
$=$ $-f^{*}{\rm Ric}(\omega^{n})+\mathscr{R}+[D_{f,\rm ram}]+2\sum_{j=1}^{q}dd^{c}(u_{D_{j}}\circ f)+2\sum_{j=1}^{q}dd^{c}\log(u_{D_{j}}\circ f)$
$=$ $-f^{*}{\rm Ric}(\omega^{n})+\mathscr{R}+[D_{f,\rm ram}]+\sum_{j=1}^{q}\left(f^{*}\omega-f^{*}[D_{j}]\right)+2\sum_{j=1}^{q}dd^{c}\log(u_{D_{j}}\circ f)$
$=$ $qf^{*}\omega-f^{*}{\rm Ric}(\omega^{n})+\mathscr{R}-f^{*}[D]+[D_{f,\rm ram}]+2\sum_{j=1}^{q}dd^{c}\log(u_{D_{j}}\circ f).$
Since $D_{j}]\leq 2dd^{c}u_{D_{j}}\leq\omega$ with $j=1,\cdots,q,$ it follows from Green-Dynkin formula and the first main theorem that
$\frac{1}{2}\int_{\partial B(r)}\log\xi d\pi_{r}-\frac{1}{2}\log\xi(o)$
$=$ $\frac{1}{4}\int_{B(r)}\Delta\log\xi dv$
$\geq$ $qT_{f}(r,\omega)+T_{f}(r,K_{N})+T(r,\mathscr{R})-N_{f}(r,D)+N(r,D_{f,\rm ram})$
$+~\frac{1}{2}\sum_{j=1}^{q}\int_{B(r)}g_{r}(o,x)\Delta\log(u_{D_{j}}\circ f)dv+O(1)$
$\geq$ $qT_{f}(r,\omega)+T_{f}(r,K_{N})+T(r,\mathscr{R})-\overline{N}_{f}(r,D)+O\big(\log T_{f}(r,\omega)\big).$
Write
$f^{*}{\rm Ric}\Psi=\frac{\sqrt{-1}}{\pi}\sum_{i,j=1}^{m}\psi_{i\bar{j}}dz_{i}\wedge d\bar{z}_{j}.$
For any $x_{0}\in M,$ one can take a normal local holomorphic coordinate $z$ near $x_{0}$ such that
$\alpha|_{x_{0}}=\frac{\sqrt{-1}}{\pi}\sum_{j=1}^{m}dz_{j}\wedge d\bar{z}_{j}.$
By Lemma 5.9
$f^{*}\Psi\wedge\alpha^{m-n}\big|_{x_{0}}$ $\leq$ $c^{-1}\left(-f^{*}{\rm Ric\Psi}\right)^{n}\wedge\alpha^{m-n}$
$=$ $\bigg(\frac{\sqrt{-1}}{\pi}\sum_{i,j=1}^{m}\psi_{i\bar{j}}dz_{i}\wedge d\bar{z}_{j}\bigg)^{n}\bigwedge\bigg(\frac{\sqrt{-1}}{\pi}\sum_{j=1}^{m}dz_{j}\wedge d\bar{z}_{j}\bigg)^{m-n}$
$=$ $c^{-1}(m-n)!\sum_{1\leq i_{1}\not=\cdots\not=i_{n}\leq m}\psi_{i_{1}\bar{i}_{1}}\cdots\psi_{i_{n}\bar{i}_{n}}\alpha^{m}$
$\leq$ $c^{-1}(m-n)!\left({\rm tr}(\psi_{i\bar{j}})\right)^{n}\alpha^{m}.$
On the other hand, we have
(24) $f^{*}{\rm Ric\Psi}\wedge\alpha^{m-1}\big|_{x_{0}}=(m-1)!{\rm tr}(\psi_{i\bar{j}})\alpha^{m},$
which leads to
$\xi^{\frac{1}{n}}\big|_{x_{0}}$ $=$ $\left(\frac{f^{*}\Psi\wedge\alpha^{m-n}}{\alpha^{m}}\right)^{\frac{1}{n}}$
$\leq$ $c^{-\frac{1}{n}}(m-n)!^{\frac{1}{n}}{\rm tr}(\psi_{i\bar{j}})$
$=$ $-\frac{c^{-\frac{1}{n}}(m-n)!^{\frac{1}{n}}}{(m-1)!}\frac{f^{*}{\rm Ric\Psi}\wedge\alpha^{m-1}}{\alpha^{m}}.$
By the arbitrariness of $x_{0},$ we show that
$\xi^{\frac{1}{n}}\leq-\frac{c^{-\frac{1}{n}}(m-n)!^{\frac{1}{n}}}{(m-1)!}\frac{f^{*}{\rm Ric\Psi}\wedge\alpha^{m-1}}{\alpha^{m}}.$
Using Corollary 5.8, it is therefore
$\int_{\partial B(r)}\log\xi d\pi_{r}\$$=$ $n\int_{\partial B(r)}\log\xi^{\frac{1}{n}}d\pi_{r}$
$\leq$ $n\log\int_{\partial B(r)}\xi^{\frac{1}{n}}d\pi_{r}\$
$\leq$ $n(1+\delta)^{2}\log\int_{B(r)}g_{r}(o,x)\xi^{\frac{1}{n}}dv+O\left(\sqrt{-\kappa(r)}r+\delta\log r\right)$
$\leq$ $n(1+\delta)^{2}\log\int_{B(r)}g_{r}(o,x)\frac{-f^{*}{\rm Ric\Psi}\wedge\alpha^{m-1}}{\alpha^{m}}dv$
$+~O\left(\sqrt{-\kappa(r)}r+\delta\log r\right)$
$=$ $n(1+\delta)^{2}\log\int_{B(r)}g_{r}(o,x)(-f^{*}{\rm Ric\Psi})\wedge\alpha^{m-1}$
$+~O\left(\sqrt{-\kappa(r)}r+\delta\log r\right)$
$=$ $n(1+\delta)^{2}\log T_{f}(r,-{\rm Ric}\Psi)+O\left(\sqrt{-\kappa(r)}r+\delta\log r\right).$
Since $N$ is compact and $\omega>0,$ we get
$T_{f}(r,-{\rm Ric}\Psi)\leq O\big(T_{f}(r,\omega)+1\big).$
This implies that
$\int_{\partial B(r)}\log\xi d\pi_{r}\leq O\left(\log T_{f}(r,\omega)+\sqrt{-\kappa(r)}r+\delta\log r\right).$
By this with (5.2), we prove the theorem. ∎
B. Ricci Non-negatively Curved Case
Assume that $M$ is non-parabolic, with non-negative Ricci curvature. Let $V(r)$ denote the Riemannian volume of geodesic ball $B(r)$ centered at $o$ with radius $r$ in $M.$ The non-parabolicity of $M$ means that
$\int_{1}^{\infty}\frac{t}{V(t)}dt<\infty.$
Thus, we have the unique minimal positive global Green function $G(o,x)$ of $\Delta/2$ for $M,$ which can be written as
$G(o,x)=2\int_{0}^{\infty}p(t,o,x)dt,$
where $p(t,o,x)$ is the heat kernel of $\Delta$ on $M.$ Let $\rho(x)$ be the Riemannian distance function of $x$ from $o.$ By Li-Yau’s estimate [10], there exist constants $A,B>0$ such that
(25) $A\int_{\rho(x)}^{\infty}\frac{t}{V(t)}dt\leq G(o,x)\leq B\int_{\rho(x)}^{\infty}\frac{t}{V(t)}dt$
holds for all $x\in M.$
5.2.1. Construction of $\Delta(r)$
For $r>0,$ define
(26) $\Delta(r)=\left\{x\in M:\ G(o,x)>A\int_{r}^{\infty}\frac{t}{V(t)}dt\right\}.$
Since $\lim_{x\to o}G(o,x)=\infty$ and $\lim_{x\to\infty}G(o,x)=0,$ we conclude that $\Delta(r)$ is a precompact domain containing $o$ satisfying that $\lim_{r\to 0}\Delta(r)\to\emptyset$ and $\lim_{r\to\infty}\Delta(r)=M.$ Moreover, the family $\{\Delta(r)\}_{r>0}$ exhausts $M.$ Hence, the boundary $\partial\Delta(r)$ of $\Delta(r)$ can be formulated as
$\partial\Delta(r)=\left\{x\in M:\ G(o,x)=A\int_{r}^{\infty}\frac{t}{V(t)}dt\right\}.$
By Sard’s theorem, $\partial\Delta(r)$ is a submanifold of $M$ for almost all $r>0.$
Set
$g_{r}(o,x)=G(o,x)-A\int_{r}^{\infty}\frac{t}{V(t)}dt.$
Note that $g_{r}(o,x)$ defines the Green function of $\Delta/2$ for $\Delta(r)$ with a pole at $o$ satisfying Dirichelet boundary condition, i.e.,
$\frac{1}{2}\Delta g_{r}(o,x)=\delta_{o}(x),\ \ \ \ ^{\forall}x\in\Delta(r);\ \ \ \ \ g_{r}(o,x)=0,\ \ \ \ ^{\forall}x\in\partial\Delta(r).$
Let $\pi_{r}$ denote the harmonic measure on $\partial\Delta(r)$ with respect to $o,$ defined by
$d\pi_{r}(x)=\frac{1}{2}\frac{\partial g_{r}(o,x)}{\partial{\vec{\nu}}}d\sigma_{r}(x),\ \ \ \ ^{\forall}x\in\partial\Delta(r),$
where $\partial/\partial\vec{\nu}$ is the inward normal derivative on $\partial\Delta(r),$ $d\sigma_{r}$ is the Riemannian area element of $\partial\Delta(r).$
Lemma 5.11.
For $x\in\partial\Delta(t),$ we have
$g_{r}(o,x)=A\int_{t}^{r}\frac{s}{V(s)}ds$
[…]
Proof.
According to the definition of Green function for $\Delta(r),$ it is immediate that for $r\geq t$
$g_{r}(o,x)$ $=$ $G(o,x)-A\int_{r}^{\infty}\frac{t}{V(t)}dt$
$=$ $G(o,x)-A\int_{t}^{\infty}\frac{s}{V(s)}ds+A\int_{t}^{r}\frac{s}{V(s)}ds$
$=$ $g_{t}(o,x)+A\int_{t}^{r}\frac{s}{V(s)}ds.$
Since $g_{t}(o,x)=0$ for $x\in\partial\Delta(t),$ we have the lemma proved. ∎
Let $\nabla$ denote the gradient operator on any Riemannian manifold. Cheng-Yau [3] proved the following lemma.
Lemma 5.12.
Let $N$ be a complete Riemannian manifold. Let $B(x_{0},r)$ be a geodesic ball centered at $x_{0}\in N$ with radius $r.$ Then, there exists a constant $c_{N}>0$ depending only on the dimension of $N$ such that
$\frac{\|\nabla u(x)\|}{u(x)}\leq\frac{c_{N}r^{2}}{r^{2}-d(x_{0},x)^{2}}\Big(|\kappa(r)|+\frac{1}{d(x_{0},x)}\Big)$
holds for any non-negative harmonic function $u$ on $B(x_{0},r),$ where $d(x_{0},x)$ is the Riemannian distance between $x_{0}$ and $x,$ and $\kappa(r)$ is the lower bound of Ricci curvature of $B(x_{0},r).$
Theorem 5.13.
There exists a constant $c_{1}>0$ such that
$\|\nabla g_{r}(o,x)\|\leq\frac{c_{1}}{r}\int_{r}^{\infty}\frac{tdt}{V(t)},\ \ \ \ ^{\forall}x\in\partial\Delta(r).$
Proof.
By the curvature assumption, ${\rm{Ric}}_{M}\geq 0.$ It follows from Lemma 5.12 and (25) that (see Remark 5 in [25] also)
$\|\nabla G(o,x)\|\leq\frac{c_{0}}{\rho(x)}G(o,x)\leq\frac{c_{0}B}{\rho(x)}\int_{\rho(x)}^{\infty}\frac{tdt}{V(t)}$
for some large constant $c_{0}>0$ which depends only on the dimension $m.$ By (25) again
$\int_{\rho(x)}^{\infty}\frac{tdt}{V(t)}\leq\int_{r}^{\infty}\frac{tdt}{V(t)},\ \ \ \ ^{\forall}x\in\partial\Delta(r),$
which gives
$\rho(x)\geq r,\ \ \ \ ^{\forall}x\in\partial\Delta(r).$
Set $c_{1}=c_{0}B,$ then we conclude that
$\|\nabla g_{r}(o,x)\|\big|_{\partial\Delta(r)}=\|\nabla G(o,x)\|\big|_{\partial\Delta(r)}\leq\frac{c_{1}}{r}\int_{r}^{\infty}\frac{tdt}{V(t)}.$
By
$\frac{\partial g_{r}(o,x)}{\partial{\vec{\nu}}}=\|\nabla g_{r}(o,x)\|,$
we obtain an estimate for the upper bound of $d\pi_{r}$ as follows.
Corollary 5.14.
There exists a constant $c_{2}>0$ such that
$d\pi_{r}\leq\frac{c_{2}}{r}\int_{r}^{\infty}\frac{tdt}{V(t)}d\sigma_{r},$
where $d\sigma_{r}$ is the Riemannian area element of $\partial\Delta(r).$
Set
(27) $H(r,\delta)=\frac{1}{r}\left(\frac{V(r)}{r}\right)^{1+\delta}\int_{r}^{\infty}\frac{tdt}{V(t)}.$
Theorem 5.15.
Let $k\geq 0$ be a locally integrable function on $M.$ Assume that $k$ is locally bounded at $o.$ Then there exists a constant $C>0$ such that for any $\delta>0,$ there exists a subset $E_{\delta}\subset(0,\infty)$ of finite Lebesgue measure such that
$\int_{\partial\Delta(r)}kd\pi_{r}\leq CH(r,\delta)\bigg(\int_{\Delta(r)}g_{r}(o,x)kdv\bigg)^{(1+\delta)^{2}}$
[…]
Proof.
Invoking Lemma 5.11, we have
$\int_{\Delta(r)}g_{r}(o,x)kdv$ $=$ $\int_{0}^{r}dt\int_{\partial\Delta(t)}g_{r}(o,x)kd\sigma_{t}$
$=$ $A\int_{0}^{r}\left(\int_{t}^{r}\frac{s}{V(s)}ds\right)dt\int_{\partial\Delta(t)}kd\sigma_{t}.$
Set
$\Lambda(r)=A\int_{0}^{r}\left(\int_{t}^{r}\frac{s}{V(s)}ds\right)dt\int_{\partial\Delta(t)}kd\sigma_{t}.$
A simple computation leads to
$\Lambda^{\prime}(r)=\frac{d\Lambda(r)}{dr}=\frac{Ar}{V(r)}\int_{0}^{r}dt\int_{\partial\Delta(t)}kd\sigma_{t}.$
In further, we have
$\frac{d}{dr}\left(\frac{V(r)\Lambda^{\prime}(r)}{r}\right)=A\int_{\partial\Delta(r)}kd\sigma_{r}.$
Apply Borel’s lemma to the left hand side of the above equality twice: one is to $V(r)\Lambda^{\prime}(r)/r$ and the other is to $\Lambda^{\prime}(r),$ then we deduce that for any $\delta>0,$ there exists a subset $E_{\delta}\subset(0,\infty)$ of finite Lebesgue measure such that
$\int_{\partial\Delta(r)}kd\sigma_{r}\leq\frac{1}{A}\left(\frac{V(r)}{r}\right)^{1+\delta}\Lambda(r)^{(1+\delta)^{2}}$
holds for all $r>0$ outside $E_{\delta}.$ Moreover, Corollary 5.14 implies that there exists a constant $c>0$ such that
$d\pi_{r}\leq\frac{c}{r}\int_{r}^{\infty}\frac{tdt}{V(t)}d\sigma_{r}.$
Set $C=c/A.$ Combining the above, we have
$\int_{\partial\Delta(r)}kd\pi_{r}$ $\leq$ $\frac{c}{Ar}\left(\frac{V(r)}{r}\right)^{1+\delta}\int_{r}^{\infty}\frac{tdt}{V(t)}\Lambda(r)^{(1+\delta)^{2}}$
$=$ $CH(r,\delta)\Lambda(r)^{(1+\delta)^{2}}.$
holds for all $r>0$ outside $E_{\delta}.$
Theorem 5.16.
Let $M$ be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let $(N,\omega)$ be a compact Kähler manifold of complex dimension not greater than that of $M.$ Let $D_{1},\cdots,D_{q}$ be effective divisors in general position on $N$ with that each $D_{j}$ is cohomologous to $\omega.$ Let $f:M\rightarrow N$ be a differentiably non-degenerate meromorphic mapping. Assume that $q\omega-{\rm Ric}(\omega^{n})>0.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$qT_{f}(r,\omega)+T_{f}(r,K_{N})+T(r,\mathscr{R})$
$\leq$ $\sum_{j=1}^{q}\overline{N}_{f}(r,D_{j})+O\big(\log T_{f}(r,\omega)+\log H(r)+\delta\log r\big)$
[…]
Proof.
Applying the similar arguments as in the proof of Theorem 5.10 and using Theorem 5.15, the theorem can be proved. ∎
6. Equidistribution of Algebroid Mappings
6.1. First Main Theorem
Let $(M,\alpha)$ be an $m$-dimensional complete non-compact Kähler manifold with Laplace-Beltrami operator $\Delta,$ and $(N,\omega)$ a Kähler projective manifold of $\dim_{\mathbb{C}}N=n\leq m.$ Let $F:M\to N$ be a $\nu$-valued algebroid mapping with $\nu$ distinct single-valued components $F_{1},\cdots,F_{\nu}.$ Set
$F^{*}\omega=\sum_{j=1}^{\nu}F_{j}^{*}\omega,\ \ \ \ F^{*}u_{D}=\sum_{j=1}^{\nu}F_{j}^{*}u_{D}.$
Given a family $\{\Delta(r)\}_{r>0}$ of precompact domains $\Delta(r)\subset M$ with smooth boundaries $\partial\Delta(r)$ which can exhaust $M.$ Fix a reference point $o\in\Delta(r).$ We use $g_{r}(o,x)$ to stand for the Green function of $\Delta/2$ for $\Delta(r)$ with a pole at $o$ satisfying Dirichlet boundary condition, and then denote by $\pi_{r}$ the harmonic measure on $\partial\Delta(r)$ with respect to $o.$
We shall define Nevanlinna’s functions of $F.$ Let $D$ be an effective divisor on $N$ cohomologous to $\omega,$ i.e., there exists a function $u_{D}\geq 0$ on $N$ such that $\omega-[D]=2dd^{c}u_{D}$ in the sense of currents. The characteristic function of $F$ with respect to $\omega$ is defined by
$T_{F}(r,\omega)=\frac{\pi^{m}}{(m-1)!\nu}\int_{\Delta(r)}g_{r}(o,x)F^{*}\omega\wedge\alpha^{m-1}.$
The proximity function of $F$ with respect to $D$ is defined by
$m_{F}(r,D)=\frac{1}{\nu}\int_{\partial\Delta(r)}F^{*}u_{D}d\pi_{r}.$
The counting function of $F$ with respect to $D$ is defined by
$N_{F}(r,D)=\frac{\pi^{m}}{(m-1)!\nu}\int_{F^{*}D\cap\Delta(r)}g_{r}(o,x)\alpha^{m-1}.$
Theorem 6.1 (First Main Theorem).
Assume that $F(o)\not\in{\rm Supp}D.$ Then
$T_{F}(r,\omega)+O(1)=m_{F}(r,D)+N_{F}(r,D).$
Proof.
Since the set of singularities of $F^{*}u_{D}$ is polar, Green-Dynkin formula can apply to each $F_{j}^{*}u_{D},$ and thus to $F^{*}u_{D}.$ Using the similar arguments as in the proof of Theorem 5.2, we can prove the theorem. ∎
6.2. Counting Function of Branch Divisors
The counting function of branch divisor $\mathscr{D}_{F}$ of $F$ is defined by
$N_{{\rm{bran}}}(r,F)=\frac{1}{\nu}N(r,\mathscr{D}_{F}),$
where
$N(r,\mathscr{D}_{F})=\frac{\pi^{m}}{(m-1)!}\int_{\mathscr{D}_{F}\cap\Delta(r)}g_{r}(o,x)\alpha^{m-1}.$
Recall that $J_{F}$ is defined by (13). Set
$N\big(r,(J_{F}=0)\big)=\frac{\pi^{m}}{(m-1)!}\int_{(J_{F}=0)\cap\Delta(r)}g_{r}(o,x)\alpha^{m-1}.$
Recall that $\jmath:N\hookrightarrow\mathbb{P}^{d}(\mathbb{C})$ is the holomorphic embedding, and $\omega_{FS}|_{N}$ is the restriction of $\omega_{FS}$ to $N.$
Theorem 6.2 (Estimate of branch divisors).
We have
$N_{\rm{bran}}(r,F)\leq(2\nu-2)\left[\frac{\omega_{FS}|_{N}}{\omega}\right]T_{F}(r,\omega)+O(1).$
Proof.
From (13), we have
$N\big(r,(J_{\psi}=0)\big)$ $=$ $\frac{1}{4}\int_{\Delta(r)}g_{r}(o,x)\Delta\log|J_{F}|^{2}dv$
$=$ $\frac{\nu-1}{2}\int_{\Delta(r)}g_{r}(o,x)\Delta\log|f_{10}\cdots f_{\nu 0}|^{2}dv$
$+\frac{1}{4}\int_{\Delta(r)}g_{r}(o,x)\Delta\log\prod_{1\leq i<j\leq\nu}\big|g_{ik}-g_{jk}\big|^{4}dv$
$=$ $\frac{\nu-1}{2}\sum_{i=1}^{\nu}\int_{\Delta(r)}g_{r}(o,x)\Delta\log|f_{i0}|^{2}dv$
$+\sum_{1\leq i<j\leq\nu}\int_{\partial\Delta(r)}\log|g_{ik}-g_{jk}|^{2}d\pi_{r}+O(1).$
Using Green-Dynkin formula, we further obtain
$N\big(r,(J_{\psi}=0)\big)$ $\leq$ $(2\nu-2)\sum_{i=1}^{\nu}N(r,g_{ik})+(2\nu-2)\sum_{i=1}^{\nu}m(r,g_{ik})+O(1)$
$=$ $(2\nu-2)\sum_{i=1}^{\nu}T(r,g_{ik})+O(1)$
$\leq$ $(2\nu-2)\sum_{i=1}^{\nu}T_{\jmath\circ F_{i}}(r,\omega_{FS})+O(1)$
$\leq$ $\nu(2\nu-2)\left[\frac{\omega_{FS}|_{N}}{\omega}\right]T_{F}(r,\omega)+O(1).$
By this, we conclude from Theorem 4.3 that
$N_{\rm{bran}}(r,F)$ $\leq$ $\frac{1}{\nu}N\big(r,(J_{\psi}=0)\big)$
$\leq$ $(2\nu-2)\left[\frac{\omega_{FS}|_{N}}{\omega}\right]T_{F}(r,\omega)+O(1).$
6.3. Second Main Theorem
Let $\mathcal{M}_{F}$ be the $\nu$-leaf complex manifold by a $\nu$-valued algebroid mapping $F:M\to N.$ Let $\pi:\mathcal{M}_{F}\to M$ be the natural projection. Equip $\mathcal{M}_{F}$ with the pullback metric $\pi^{*}g$ induced by $g$ on $M.$ Note that $\pi^{*}g$ is a positive semi-definite Kähler metric on $\mathcal{M}_{F},$ i.e., $\pi^{*}g$ is a Kähler metric on $\mathcal{M}_{F}\setminus\pi^{-1}(\mathscr{B}_{F})$ and degenerate on $\pi^{-1}(\mathscr{B}_{F}).$ Since $\mathscr{B}_{F}$ is a polar set of ${\rm codim}_{\mathbb{C}}\mathscr{B}_{F}=1,$ the Green-Dynkin formula still works on each leaf of $(\mathcal{M}_{F},\pi^{*}g)$ for the following reasons. Set $K=\pi^{-1}(\mathscr{B}_{F}).$ Take a sequence $\{U_{n}(K)\}_{n=1}^{\infty}$ of neighborhoods $U_{n}(K)$ of $K$ such that
$\cdots\subseteq\overline{U_{k+1}(K)}\subseteq U_{k}(K)\subseteq\cdots\subseteq\overline{U_{2}(K)}\subseteq U_{1}(K),\ \ \ \ \bigcap_{n\geq 1}U_{n}(K)=K.$
Let $\{\tau_{n},\theta_{n}\}$ be a partition of the unity subordinate to the open covering $\{\mathcal{M}_{F}\setminus U_{n+1}(K),U_{n}(K)\}$ of $\mathcal{M}_{F}.$ So, $0\leq\tau_{n}\leq 1,0\leq\theta_{n}\leq 1$ and $\tau_{n}+\theta_{n}=1.$ Fix a Kähler metric $g_{0}$ on $\mathcal{M}_{F}$ with Kähler form $\omega_{0}.$ Set $g_{n}=\pi^{*}g+n^{-2}\theta_{n}g_{0}.$ Since $g_{0}$ is positive definite on $\mathcal{M}_{F},$ we see that $g_{n}$ is also positive definite on $\mathcal{M}_{F}.$ Moreover, since $d(\pi^{*}\omega+n^{-2}\theta_{n}\omega_{0})=\pi^{*}d\omega+n^{-2}\theta_{n}d\omega_{0}=0,$ we conclude that $g_{n}$ is a Kähler metric on $\mathcal{M}_{F}$ with $g_{n}=\pi^{*}g$ on $\mathcal{M}_{F}\setminus U_{n}(K).$ Let $\tilde{m},\tilde{m}_{n}$ stand for the Riemannian measures of $\pi^{*}g,g_{n},$ respectively. Since ${\rm codim}_{\mathbb{C}}\mathscr{B}_{F}=1$ and $g_{n}\to\pi^{*}g$ as $n\to+\infty,$ we deduce that $\tilde{m}_{n}\to\tilde{m}$ weakly as $n\to+\infty.$ Let $\tilde{\Delta},\tilde{\Delta}_{n}$ stand for the Laplace-Beltrami operators of $\pi^{*}g,g_{n},$ respectively, in which $\tilde{\Delta}$ is well-defined on $\mathcal{M}_{F}\setminus K.$ According to $g_{n}\to\pi^{*}g$ and $\tilde{m}_{n}\to\tilde{m}$ weakly as $n\to+\infty,$ we obtain $\tilde{\Delta}_{n}\to\tilde{\Delta}$ weakly as $n\to+\infty.$ Taking a precompact domain $\tilde{\Omega}$ with smooth boundary $\partial\tilde{\Omega}$ in a leaf of $\mathcal{M}_{F}.$ We fix a reference point $\tilde{o}\in\tilde{\Omega}$ such that $\tilde{o}\not\in K.$ Denote by $\tilde{g}_{\tilde{\Omega}}(\tilde{o},x),\tilde{g}_{\tilde{\Omega},n}(\tilde{o},x)$ the Green functions of $\tilde{\Delta}/2,\tilde{\Delta}_{n}/2$ for $\tilde{\Omega}$ with a pole at $\tilde{o}$ satisfying Dirichlet boundary condition, respectively, in which $\tilde{g}_{\tilde{\Omega}}(\tilde{o},x)$ is well-defined on $\overline{\tilde{\Omega}}\setminus K.$ Again, we denote by $\tilde{\pi}_{\partial\tilde{\Omega}},\tilde{\pi}_{\partial\tilde{\Omega},n}$ the harmonic measures of $\pi^{*}g,g_{n}$ on $\partial\tilde{\Omega}$ with respect to $\tilde{o}$ respectively, in which $\tilde{\pi}_{\partial\tilde{\Omega}}$ is well-defined on $\partial{\tilde{\Omega}}\setminus K.$ Since $\mathscr{B}_{F}$ is a polar set, the previous arguments can follow that $\tilde{\Delta}_{n}\tilde{g}_{\tilde{\Omega},n}(\tilde{o},x)\to\tilde{\Delta}\tilde{g}_{\tilde{\Omega}}(\tilde{o},x)$ in the sense of distributions and $\tilde{\pi}_{\partial\tilde{\Omega},n}\to\tilde{\pi}_{\partial\tilde{\Omega}}$ weakly as $n\to+\infty.$ Applying Green-Dynkin formula, we get
$\int_{\partial\tilde{\Omega}}u(x)d\tilde{\pi}_{\partial\tilde{\Omega},n}(x)-u(\tilde{o})=\frac{1}{2}\int_{\tilde{\Omega}}\tilde{g}_{\tilde{\Omega},n}(\tilde{o},x)\tilde{\Delta}_{n}u(x)d\tilde{v}_{n}(x),\ \ \ \ ^{\forall}n\geq 1$
[…]
$\int_{\partial\tilde{\Omega}}u(x)d\tilde{\pi}_{\partial\tilde{\Omega}}(x)-u(\tilde{o})=\frac{1}{2}\int_{\tilde{\Omega}}\tilde{g}_{\tilde{\Omega}}(\tilde{o},x)\tilde{\Delta}u(x)d\tilde{v}(x).$
It implies that Dynkin formula works on each leaf of $\mathcal{M}_{F}$ under the pullback metric $\pi^{*}g.$ According to Corollary 3.6, $F$ can lift to a meromorphic mapping $f:\mathcal{M}_{F}\to N$ such that $F=f\circ\pi^{-1}.$ In further, it is not difficult to examine that Theorem 5.2, Theorem 5.10 and Theorem 5.16 also apply to $f$ on each leaf of $\mathcal{M}_{F}$ under the corresponding conditions.
A. Non-positively Curved Case
With the pullback metric $\pi^{*}g,$ we see that $\mathcal{M}_{F}$ has non-positive sectional curvature outside $\pi^{-1}(\mathscr{B}_{F}).$ Let $\tilde{\alpha}=\pi^{*}\alpha$ and $\tilde{\mathscr{R}}=\pi^{*}{\mathscr{R}}$ be the Kähler form and Ricci form of $\pi^{*}g,$ respectively, which are well-defined on $\mathcal{M}_{F}\setminus\pi^{-1}(\mathscr{B}_{F}).$ Given an effective divisor $D$ cohomologous to $\omega$ with associated function $u_{D}.$ Review the Nenvanlinna’s functions of $F$ by setting $\Delta(r)=B(r)$ in Section 6.1.
Let $\tilde{B}_{1}(r),\cdots,\tilde{B}_{\nu}(r)$ stand for the $\nu$ connected components of $\pi^{-1}(B(r))$ centered at $\tilde{o}_{1},\cdots,\tilde{o}_{\nu}$ with that $\pi(\tilde{o}_{k})=o$ for each $k,$ respectively. Define
$\tilde{g}_{r}(\tilde{o}_{k},x)=g_{r}(o,\pi(x)),\ \ \ \ d\tilde{\pi}_{k,r}(x)=d\pi_{r}(\pi(x)).$
Note that outside $\pi^{-1}(\mathscr{B}_{F}),$ $\tilde{g}_{r}(\tilde{o}_{k},x)$ is the Green function of $\tilde{\Delta}/2$ for $\tilde{B}_{k}(r)$ with a pole at $\tilde{o}_{k}$ satisfying Dirichlet boundary condition, and moreover, $\tilde{\pi}_{k,r}$ is the harmonic measure on $\partial{\tilde{B}}_{k}(r)$ with respect to $\tilde{o}_{k}.$ For each $k,$ put
$T_{f,k}(r,\omega)$ $=$ $\frac{\pi^{m}}{(m-1)!}\int_{\tilde{B}_{k}(r)}\tilde{g}_{r}(\tilde{o}_{k},x)f^{*}\omega\wedge\tilde{\alpha}^{m-1},$
$m_{f,k}(r,D)$ $=$ $\int_{\partial{\tilde{B}}_{k}(r)}f^{*}u_{D}d\tilde{\pi}_{r},$
$N_{f,k}(r,D)$ $=$ $\frac{\pi^{m}}{(m-1)!}\int_{f^{*}(D)\cap\tilde{B}_{k}(r)}\tilde{g}_{r}(\tilde{o}_{k},x)\tilde{\alpha}^{m-1},$
$\overline{N}_{f,k}(r,D)$ $=$ $\frac{\pi^{m}}{(m-1)!}\int_{f^{-1}(D)\cap\tilde{B}_{k}(r)}\tilde{g}_{r}(\tilde{o}_{k},x)\tilde{\alpha}^{m-1}$
and
$T_{k}(r,\tilde{\mathscr{R}})=\frac{\pi^{m}}{(m-1)!}\int_{\tilde{B}_{k}(r)}\tilde{g}_{r}(\tilde{o}_{k},x)\tilde{\mathscr{R}}\wedge\tilde{\alpha}^{m-1}.$
When $F$ is differentiably non-degenerate, i.e., at least one single-valued component of $F$ is differentiably non-degenerate, we can yield from Theorem 5.10 that for effective divisors $D_{1},\cdots,D_{q}$ in general position on $N$ such that each $D_{j}$ is cohomologous to $\omega,$ there is a subset $E_{\delta}\subset(0,\infty)$ of finite Lebesgue measure such that for $k=1,\cdots,\nu$
[…]
$\leq$ $\sum_{j=1}^{q}\overline{N}_{f,k}(r,D_{j})+O\big(\log T_{f,k}(r,\omega)+\sqrt{-\kappa_{k}(r)}r+\delta\log r\big)$
[…]
Theorem 6.3 (Second Main Theorem).
Let $M$ be a complete non-compact Kähler manifold with non-positive sectional curvature. Let $(N,\omega)$ be a Kähler projective manifold of dimension not greater than that of $M.$ Let $D_{1},\cdots,D_{q}$ be effective divisors in general position on $N$ such that each $D_{j}$ is cohomologous to $\omega.$ Let $F:M\rightarrow N$ be a differentiably non-degenerate $\nu$-valued algebroid mapping. Assume that $q\omega-{\rm Ric}(\omega^{n})>0.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$\left(q-(2\nu-2)\left[\frac{\omega_{FS}|_{N}}{\omega}\right]\right)T_{F}(r,\omega)+T_{F}(r,K_{N})+T(r,\mathscr{R})$
$\leq$ $\sum_{j=1}^{q}\overline{N}_{F}(r,D_{j})+O\big(\log T_{F}(r,\omega)+\sqrt{-\kappa(r)}r+\delta\log r\big)$
[…]
Proof.
It is noticed that $\pi:\mathcal{M}_{F}\setminus{\pi^{-1}(\mathscr{B}_{F})}\rightarrow M\setminus{\mathscr{B}_{F}}$ is a $\nu$-sheeted analytic covering, and two distinct connected components $\tilde{B}_{i}(r),\tilde{B}_{j}(r)$ of $\pi^{-1}(B(r))$ meet only at points in $\pi^{-1}(\mathscr{B}_{F}).$ In addition, any point in $F^{-1}(D)$ is possibly a branch point of $w,$ or any point in $f^{-1}(D)$ is possibly a ramification point of $\pi,$ where $D=D_{1}+\cdots+D_{q}.$ Hence, it yields from $F=f\circ\pi^{-1}$ that
$\frac{1}{\nu}\sum_{k=1}^{\nu}\sum_{j=1}^{q}\overline{N}_{f,k}(r,D_{j})\leq\sum_{j=1}^{q}\overline{N}_{F}(r,D_{j})+N_{\rm{bran}}(r,F).$
[…]
$T_{F}(r,\omega)\leq\frac{1}{\nu}\sum_{k=1}^{\nu}T_{f,k}(r,\omega)\leq T_{F}(r,\omega)+N_{\rm{bran}}(r,F).$
Moreover, since the curvature function on $M$ is continuous and the curvature of $\pi^{*}g$ outside $\pi^{-1}(\mathscr{B}_{F})$ matches the curvature of $g$ outside $\mathscr{B}_{F}$ locally, one obtains $\kappa_{k}=\kappa$ for $k=1,\cdots,\nu.$ Put together the above with (6.3), then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(0,\infty)$ of finite Lebesgue measure such that
$qT_{F}(r,\omega)+T_{F}(r,K_{N})+T(r,\mathscr{R})$
[…]
$\leq$ $\frac{1}{\nu}\sum_{k=1}^{\nu}\bigg(\sum_{j=1}^{q}\overline{N}_{f,k}(r,D_{j})+O\big(\log T_{f,k}(r,\omega)+\sqrt{-\kappa_{k}(r)}r+\delta\log r\big)\bigg)$
$\leq$ $\sum_{j=1}^{q}\overline{N}_{F}(r,D_{j})+N_{\rm{bran}}(r,F)$
$+~O\Big(\log\big(T_{F}(r,\omega)+N_{\rm{bran}}(r,F)\big)+\sqrt{-\kappa(r)}r+\delta\log r\Big)$
holds for all $r>0$ outside $E_{\delta}.$ Using Theorem 6.2, then we have the theorem proved. ∎
Let $\mathscr{O}(1)$ stand for the hyperplane line bundle over $\mathbb{P}^{n}(\mathbb{C}).$ It yields from $c_{1}(K_{\mathbb{P}^{n}(\mathbb{C})})=-(n+1)c_{1}(\mathscr{O}(1))$ that
Corollary 6.4.
Let $M$ be a complete non-compact Kähler manifold with non-positive sectional curvature. Let $H_{1},\cdots,H_{q}$ be hyperplanes of $\mathbb{P}^{n}(\mathbb{C})$ in general position. Let $F:M\rightarrow\mathbb{P}^{n}(\mathbb{C})$ be a differentiably non-degenerate $\nu$-valued algebroid mapping. Assume that $m\geq n.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$\left(q-2\nu-n+1\right)T_{F}(r,\omega_{FS})+T(r,\mathscr{R})$
$\leq$ $\sum_{j=1}^{q}\overline{N}_{F}(r,H_{j})+O\big(\log T_{F}(r,\omega_{FS})+\sqrt{-\kappa(r)}r+\delta\log r\big)$
[…]
B. Ricci Non-negatively Curved Case
Assume that $M$ is non-parabolic. Under the pullback metric $\pi^{*}g,$ one can see that $\mathcal{M}_{F}$ is also non-parabolic, with non-negative Ricci curvature outside $\pi^{-1}(\mathscr{B}_{F}).$ Similarly, we can have $T_{F}(r,\omega),m_{F}(r,D),N_{F}(r,D),\overline{N}_{F}(r,D)$ and $T(r,\mathscr{R})$ on $\Delta(r)$ defined by (26) in Section 5.2.1. By the similar arguments as in the proof of Theorem 6.3, we also conclude from Theorem 5.16 that
Theorem 6.5 (Second Main Theorem).
Let $M$ be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let $(N,\omega)$ be a Kähler projective manifold of dimension not greater than that of $M.$ Let $D_{1},\cdots,D_{q}$ be effective divisors in general position on $N$ such that each $D_{j}$ is cohomologous to $\omega.$ Let $F:M\rightarrow N$ be a differentiably non-degenerate $\nu$-valued algebroid mapping. Assume that $q\omega-{\rm Ric}(\omega^{n})>0.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$\left(q-(2\nu-2)\left[\frac{\omega_{FS}|_{N}}{\omega}\right]\right)T_{F}(r,\omega)+T_{F}(r,K_{N})+T(r,\mathscr{R})$
$\leq$ $\sum_{j=1}^{q}\overline{N}_{F}(r,D_{j})+O\big(\log T_{F}(r,\omega)+\log H(r)+\delta\log r\big)$
[…]
Similarly, we have:
Corollary 6.6.
Let $M$ be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Let $H_{1},\cdots,H_{q}$ be hyperplanes of $\mathbb{P}^{n}(\mathbb{C})$ in general position. Let $F:M\rightarrow\mathbb{P}^{n}(\mathbb{C})$ be a differentiably non-degenerate $\nu$-valued algebroid mapping. Assume that $m\geq n.$ Then for any $\delta>0,$ there exists a subset $E_{\delta}\subset(1,\infty)$ of finite Lebesgue measure such that
$\left(q-2\nu-n+1\right)T_{F}(r,\omega_{FS})+T(r,\mathscr{R})$
$\leq$ $\sum_{j=1}^{q}\overline{N}_{F}(r,H_{j})+O\big(\log T_{F}(r,\omega_{FS})+\log H(r)+\delta\log r\big)$
[…]